### v1

# Neuro-Symbolic Generation of Explanations for Robot Policies with Weighted Signal Temporal Logic

Abstract
Neural network-based policies have demonstrated success in many robotic applications, but often lack human-explanability, which poses challenges in safety-critical deployments. To address this, we propose a neuro-symbolic explanation framework that generates a weighted signal temporal logic (wSTL) specification to describe a robot policy in a interpretable form. Existing methods typically produce explanations that are verbose and inconsistent, which hinders explainability, and loose, which do not give meaningful insights into the underlying policy. We address these issues by introducing a simplification process consisting of predicate filtering, regularization, and iterative pruning. We also introduce three novel explainability evaluation metrics—conciseness, consistency, and strictness—to assess explanation quality beyond conventional classification metrics. Our method is validated in three simulated robotic environments, where it outperforms baselines in generating concise, consistent, and strict wSTL explanations without sacrificing classification accuracy. This work bridges neural network policy learning with formal methods, contributing to safer and more transparent decision-making in robotics.
I Introduction
Recent innovations in learning-based methods, particularly those using neural networks, have significantly improved both high-level decision-making and low-level control in robotics. However, these neural network policies are inherently black-box in nature, which poses significant concerns for safety-critical robotic applications, such as autonomous vehicles, where explainability is paramount. To address these concerns, various approaches have been explored to enhance explainability of learned policies [1, 2, 3], including post-hoc analysis techniques and methods to learn policies that are inherently explainable.
In this paper, we focus on post-hoc analysis to enable explanation of any given robot policy. While existing approaches, like saliency maps [4] and feature attribution [5], provide valuable information, they often lack language-based explanations, which limits human-interpretability. Transformer-based language models have recently shown promising results for providing language-based explanations [6, 7, 8], but they do not support formal verification of the output explanation, which limits their applicability in domains requiring strong safety and correctness assurances [9].
Formal methods offer an alternative that emphasizes verifiability. In particular, temporal logic has been used to specify robotic tasks and system constraints in an interpretable yet mathematically precise manner that supports verification [10, 11, 12]. Recent efforts have also used temporal logic, primarily linear temporal logic (LTL) and signal temporal logic (STL), to infer explanations of given trajectory data. For example, [13, 14, 15] use a decision tree classifier to infer the LTL or STL specification that best explains the difference between two classes of trajectories. [16] uses a greedy explanation search method over explanation candidates in LTL driven by policy similarity metrics. However, these methods are limited in their ability to infer specifications where certain predicates are more critical than others. For instance, [15] can infer such relations only for conjunctions. Weighted signal temporal logic (wSTL) [17] extends LTL and STL by introducing importance weights over predicates, which can highlight the relative priority of various task components. The expressiveness of wSTL thus makes it particularly suitable for describing robotic behaviors in settings with competing or hierarchical objectives.
Figure 1: Existing methods for neuro-symbolic post-hoc explainability produce specifications that are verbose (i.e., include unnecessary predicates), inconsistent across training seeds, and loose (e.g., “catch-all” specifications that do not provide useful insights). Our proposed network, TLNet, addresses these concerns.
Figure 2: We propose a neuro-symbolic approach for generating an explanation of a given robot policy using wSTL. Our network architecture, TLNet, includes a simplification process that ensures resulting explanations balance classification accuracy with conciseness, consistency, and strictness.
The presence of real-valued weights in wSTL has enabled the use of neural network approaches for inferring wSTL specifications that explain given trajectory data. These methods have shown promising results, where they show high classification performance for various applications, including anomaly detection in naval surveillance [18], fault diagnosis [19], and weather forecasting [20]. However, existing work only focuses on classification accuracy, which does not fully capture the actual explanability of learned specifications. For example, as shown in Figure 1, these methods often produce specifications that are unnecessarily verbose (i.e., they include many unnecessary predicates), inconsistent across seeds, and/or loose, such that they are satisfied but do not provide useful insights (e.g., “catch-all” specifications that include many conjunctions or overly simplified specifications that do not fully capture the task or constraints). Some works address verbosity by incorporating heuristic simplification methods, like only keeping the top-$k$ weighted predicates [18, 20] or using a greedy process to remove weights [21]. However, as we show in our empirical results, no existing methods effectively balance classification accuracy with verbosity, inconsistency, and looseness.
In this work, we propose TLNet, a neural network architecture for inferring wSTL explanations of robot policies that are concise (i.e., include as few predicates as needed), consistent (i.e., produce similar explanations across training seeds), and strict (i.e., contain as few disjunctions as needed). Our method targets a subclass of wSTL structures with arbitrary conjunctions and disjunctions of predicates in temporal clauses. TLNet includes a multi-step simplification process that reduces specification verbosity and improves consistency, while ensuring strictness is maintained, as shown in Figure 2. This process includes a predicate filtering step, a regularized loss function, and iterative weight pruning. We also provide theoretical justification for our wSTL-based neural network architecture, which essentially reduce learning parameters by the distributivity of importance weight, and introduce novel metrics for quantifying explainability—conciseness, consistency, and strictness—that complement standard classification accuracy, addressing a significant gap in current evaluation practices. We demonstrate our method in three simulated robotics environments that vary in continuity, dimensionality, task structure, and policy complexity, and show that it generally outperforms baselines with respect to classification accuracy and our proposed explainability metrics. TLNet bridges the gap between neural network-based policy learning and providing formal explanations of learned policies, offering a step toward safer and more transparent robotic systems.
II Preliminaries
Let a trajectory (or a signal) of $n$-dimensional states be a discrete-time series $\tau=s_{0},s_{1},..,s_{H}$, where $s_{t}\in S$ is a state at timestep $t\in\mathbb{Z}_{>0}$, $S\subseteq\mathbb{R}^{n}$ is the state space, and $H$ is the time horizon. We also define an interval $I$ as $I:=[a,b]=\{k\in\mathbb{Z}_{\geq 0}|a\leq k\leq b;a,b\in\mathbb{Z}_{\geq 0}\}$, $\lvert I\rvert$ as the cardinality of $I$, and $t+I$ as the interval $[t+a,t+b]$.
II-A Weighted Signal Temporal Logic
We consider wSTL syntax [17] defined recursively as $\phi:=\top\mid\psi\mid\neg\phi\mid\bigwedge_{i=1}^{N}{}^{w}\phi_{i}\mid\bigvee_{i=1}^{N}{}^{w}\phi_{i}\mid\mathcal{G}_{I}^{w}\phi\mid\mathcal{F}_{I}^{w}\phi$, where $\top$ is the logical True value, $\psi:=f(s)\geq c$ is an atomic predicate given a real-valued function $f:S\rightarrow\mathbb{R}$ over $S$ with constant $c\in\mathbb{R}$, and $\neg$, $\wedge$, and $\vee$ are Boolean negation, conjunction, and disjunction operators, respectively. For conjunctive and disjunctive operators, we define the weights as $w:=[w_{i}]_{i=1:N}\in\mathbb{R}^{N}_{>0}$, where $N$ is the number of subformulae in the operator. For temporal operators, we define the weights as $w:=[w_{k}]_{k\in I}\in\mathbb{R}^{\lvert I\rvert}_{>0}$ for interval $I$. For notational simplicity, we use the following notation interchangeably: $\bigwedge_{i=1}^{N}{}^{w}\phi_{i}=\bigwedge_{i=1}^{N}w_{i}\phi_{i}$ and $\bigvee_{i=1}^{N}{}^{w}\phi_{i}=\bigvee_{i=1}^{N}w_{i}\phi_{i}$.
II-B Normalized wSTL
Our neural network architecture proposed in Section III-A requires the weights within each operator to be normalized. We discuss the reasons for this requirement in Section III-E. Given this requirement, we define normalized wSTL syntax recursively as $\phi:=\top\mid\psi\mid\neg\phi\mid\bigwedge_{i=1}^{N}{}^{w}\phi_{i}\mid\bigvee_{i=1}^{N}{}^{w}\phi_{i}\mid\mathcal{G}_{I}^{w}\phi\mid\mathcal{F}_{I}^{w}\phi_{i},$ where all terminology carries over from wSTL, except we further require $\sum_{i=1}^{N}w_{i}=1$ and $\sum_{k\in I}w_{k}=1$, where $0\leq w_{i},w_{k}\leq 1$. We allow the weights to be zero because the zero weights do not influence the overall robustness [18].
To avoid large input values to our neural network, we also define normalized quantitative satisfaction semantics, or robustness, by extending [17] as follows,
$r(\psi,\tau,t)$ ::=:= $\begin{cases}\frac{f(s_{t})-c}{\sup_{s\in S}f(s)-c}&\text{if $f(s_{t})-c\geq 0$,}\\ \frac{f(s_{t})-c}{\inf_{s\in S}f(s_{t})-c}&\text{otherwise}\end{cases},$ (1)
$r(\neg\phi,\tau,t)$ ::=:= $-r(\phi,\tau,t),$ (2)
$r\left(\bigwedge_{i=1}^{N}{}^{w}\phi_{i},\tau,t\right)$ ::=:= $\otimes^{\wedge}(w,[r(\phi_{i},\tau,t)]_{i=1:N}),$ (3)
$r\left(\bigvee_{i=1}^{N}{}^{w}\phi_{i},\tau,t\right)$ ::=:= $\oplus^{\vee}(w,[r(\phi_{i},\tau,t)]_{i=1:N}).$ (4)
$r\left(\mathcal{G}^{w}_{I}\phi,\tau,t\right)$ ::=:= $\otimes^{\mathcal{G}}_{I}(w,[r(\phi,\tau,t^{\prime})]_{t^{\prime}\in t+I}),$ (5)
$r\left(\mathcal{F}^{w}_{I}\phi,\tau,t\right)$ ::=:= $\oplus^{\mathcal{F}}_{I}(w,[r(\phi,\tau,t^{\prime})]_{t^{\prime}\in t+I}),$ (6)
where the aggregation functions $\otimes^{\wedge}:\mathbb{R}^{N}_{\geq 0}\times\mathbb{R}^{N}\rightarrow\mathbb{R},\otimes^{\vee}:\mathbb{R}^{N}_{\geq 0}\times\mathbb{R}^{N}\rightarrow\mathbb{R},\otimes^{\mathcal{G}}:\mathbb{R}^{\lvert I\rvert}_{\geq 0}\times\mathbb{R}^{\lvert I\rvert}\rightarrow\mathbb{R}$, and $\otimes^{\mathcal{F}}:\mathbb{R}^{\lvert I\rvert}_{\geq 0}\times\mathbb{R}^{\lvert I\rvert}\rightarrow\mathbb{R}$ correspond to the operators $\wedge,\vee,\mathcal{G}$, and $\mathcal{F}$, respectively. For notational simplicity, we interchangeably use $r(\phi,\tau,t)$ and $r_{\tau t}^{\phi}$.
Building from [18], we define the aggregation functions as follows,
$\otimes^{\wedge}(w,[r_{\tau t}^{\phi_{i}}]_{i=1:N})$ $:=\frac{\sum_{i=1}^{N}w_{i}r_{\tau t}^{\phi_{i}}\exp(\frac{-r_{\tau t}^{\phi_{i}}}{\sigma})}{\sum_{i=1}^{N}w_{i}\exp(\frac{-r_{\tau t}^{\phi_{i}}}{\sigma})},$ (7)
$\oplus^{\vee}(w,[r_{\tau t}^{\phi_{i}}]_{i=1:N})$ $:=\frac{\sum_{i=1}^{N}w_{i}r_{\tau t}^{\phi_{i}}\exp(\frac{r_{\tau t}^{\phi_{i}}}{\sigma})}{\sum_{i=1}^{N}w_{i}\exp(\frac{r_{\tau t}^{\phi_{i}}}{\sigma})},$ (8)
$\otimes^{\mathcal{G}}[w,[r_{\tau t^{\prime}}^{\phi}]_{t^{\prime}\in t+I})$ $:=\frac{\sum_{t^{\prime}\in t+I}w_{t^{\prime}}r_{\tau t^{\prime}}^{\phi}\exp(\frac{-r_{\tau t^{\prime}}^{\phi}}{\sigma})}{\sum_{t^{\prime}\in t+I}w_{t^{\prime}}\exp(\frac{-r_{\tau t^{\prime}}^{\phi}}{\sigma})},$ (9)
$\oplus^{\mathcal{F}}[w,[r_{\tau t^{\prime}}^{\phi}]_{t^{\prime}\in t+I})$ $:=\frac{\sum_{t^{\prime}\in t+I}w_{t^{\prime}}r_{\tau t^{\prime}}^{\phi}\exp(\frac{r_{\tau t^{\prime}}^{\phi}}{\sigma})}{\sum_{t^{\prime}\in t+I}w_{t^{\prime}}\exp(\frac{r_{\tau t^{\prime}}^{\phi}}{\sigma})},$ (10)
where $\sigma>0$ is a scaling constant.
II-C Problem Statement
We consider the following binary classification task. Let $\mathcal{T}=\Set{(\tau_{i},l_{i})}{i=1,\dots,N_{\mathrm{data}}}$ be a labeled dataset, where $\tau_{i}\in\mathbb{R}^{n\times H}$ is the $i^{th}$ trajectory with label $l_{i}\in\{1,-1\}$ and $N_{\mathrm{data}}\in\mathbb{Z}_{>0}$ is the number of data points in the dataset.
Problem 1.
Given a dataset $\mathcal{T}$, find a wSTL specification $\phi$ that balances classification accuracy with conciseness, consistency, and strictness.
III Methodology
We first introduce our neural network architecture, TLNet, which allows us to represent wSTL structures in a fully differentiable form. We then introduce our simplification process, which enhances the explainability of the generated explanations, and our trajectory sampling process. Finally, we propose novel evaluation metrics for explanability—conciseness, consistency, and strictness—and provide theoretical justification for our network architecture. Algorithm 1 summarizes our overall approach.
Algorithm 1 Explanation Generation
1:Dataset $\mathcal{T}$, predicates $\Psi=\Set{\psi_{i}}{i=1,\dots,N_{\mathrm{AP}}}$, similarity threshold $S_{\mathrm{th}}$, number of pruning iterations $N_{\mathrm{pr}}$, number of weights to prune $N_{\mathbf{w}}$, loss function $\tilde{\mathcal{L}}.$
2:Explanation $\phi$
3:Evaluate robustness of trajectories in $\mathcal{T}$ using $\Psi$.
4:Filter out predicates whose similarities are above $S_{\mathrm{th}}$.
5:Initialize the neural network $f_{\mathbf{w}}$ with parameters $\mathbf{w}$.
6:Optimize $\mathbf{w}$ to minimize $\tilde{\mathcal{L}}$.
7:for $i=1:N_{\mathrm{pr}}$ do
8:     Prune zeroed and smallest $N_{\mathbf{w}}$ parameters in $\mathbf{w}$.
9:     Re-normalize $\mathbf{w}$.
10:     Optimize $\mathbf{w}$ to minimize $\tilde{\mathcal{L}}$.
11:end for
12:return $\phi$ generated from $\mathbf{w}$ and $\Psi$ using Equation 11.
III-A Neural Network Architecture
Let $\Psi=\Set{\psi_{i}}{i=1,\dots,N_{\mathrm{AP}}}$ be a predefined set of atomic predicates and $\Psi_{\neg}=\Set{\neg\psi_{i}}{i=1,\dots,N_{\mathrm{AP}}}$ be a set of their negations. We focus on representing wSTL structures composed of temporal operators containing these predicates in conjunctive normal form (CNF), as follows,
$\phi=0.5\mathcal{F}_{I}^{w^{U}}\left[\bigwedge_{i=1}^{N_{\mathrm{AP}}}\bigvee_{j=1}^{2N_{\mathrm{AP}}}w_{ij}^{\mathcal{F}}\psi_{j}\right]\\ \wedge 0.5\mathcal{G}_{I}^{w^{U}}\left[\bigwedge_{i=1}^{N_{\mathrm{AP}}}\bigvee_{j=1}^{2N_{\mathrm{AP}}}w_{ij}^{\mathcal{G}}\psi_{j}\right],$ (11)
where $w_{ij}^{\mathcal{F}},w_{ij}^{\mathcal{G}}\in[0,1]$ are importance weights and $\psi_{j}\in\Psi\cup\Psi_{\neg}$ are predefined atomic predicates defining specification $\phi$. We choose this class of structures because they allow us to represent a general class of robot policies aiming to complete a task while satisfying a set of constraints [22, 16]. We define the weights $w^{U}$ for temporal operators over interval $I=[0,H]$ to be $w^{U}=[w^{U}_{k}=1/\lvert I\rvert]_{k=1:\lvert I\rvert}$ for trajectory $\tau$. We define the conjunction weights of the temporal clauses to be 0.5 to equally weigh the importance of the task ($\mathcal{F}$) and constraint ($\mathcal{G}$) parts of the explanation.
The architecture of TLNet is designed to ensure the built-in architectural interpretability of the network weights, achieved by the bijective correspondence of the network weights and the importance weights of Equation 11. In Section III-E, we detail the theoretical justification of the formula manipulation to aggregate the weights of the conjunctive and disjunctive clauses in Equation 11. The output of TLNet is the robustness of the input trajectory with respect to the specification defined by the current network weights.
An example TLNet architecture for a dataset containing two predicates is shown in Figure 3. The specification represented by this architecture is $\phi=0.5\mathcal{F}_{I}^{w^{U}}[(w_{11}^{\mathcal{F}}\psi_{1}\vee w_{12}^{\mathcal{F}}\psi_{2}\vee w_{13}^{\mathcal{F}}\neg\psi_{1}\vee w_{14}^{\mathcal{F}}\neg\psi_{2})\wedge(w_{21}^{\mathcal{F}}\psi_{1}\vee w_{22}^{\mathcal{F}}\psi_{2}\vee w_{23}^{\mathcal{F}}\neg\psi_{1}\vee w_{24}^{\mathcal{F}}\neg\psi_{2})]\wedge 0.5\mathcal{G}_{I}^{w^{U}}[(w_{11}^{\mathcal{G}}\psi_{1}\vee w_{12}^{\mathcal{G}}\psi_{2}\vee w_{13}^{\mathcal{G}}\neg\psi_{1}\vee w_{14}^{\mathcal{G}}\neg\psi_{2})\wedge(w_{21}^{\mathcal{G}}\psi_{1}\vee w_{22}^{\mathcal{G}}\psi_{2}\vee w_{23}^{\mathcal{G}}\neg\psi_{1}\vee w_{24}^{\mathcal{G}}\neg\psi_{2})]$.
Figure 3: Example architecture of the proposed neural network when there are two predicates to consider ($\psi_{1}$ and $\psi_{2}$). $\mathbf{w}^{\mathcal{F}}$ and $\mathbf{w}^{\mathcal{G}}$ are weight vectors that are optimized. $\phi$ is the inferred explanation to the given trajectory.
III-B Simplification Process
A key component of TLNet is its simplification process, which aims to produce more human-interpretable wSTL explanations while preserving classification performance. This process consists of three steps: a predicate filter, loss regularization, and weight pruning.
The predicate filter eliminates predicates that exhibit similar behavior across both positive and negative trajectory sets, as they contribute little to discriminating between those sets and may introduce unnecessary complexity. Given a dataset of trajectories $\mathcal{D}$ and a predicate $\psi\in\Psi$, we define the robustness distribution as
$d(\mathcal{D},\psi)=\frac{1}{|\mathcal{D}|}\left[\lvert\mathcal{D}_{1}\rvert,\lvert\mathcal{D}_{2}\rvert,\lvert\mathcal{D}_{3}\rvert\right],$ (12)
where
$\mathcal{D}_{1}$ :=$:=\set{\tau}{\tau\in\mathcal{D},r(\psi,\tau,t^{\prime})\geq 0,\forall t^{\prime}\in t+I},$ (13)
$\mathcal{D}_{2}$ $:=\mathcal{D}\setminus(\mathcal{D}_{1}\cup\mathcal{D}_{3}),$ (14)
$\mathcal{D}_{3}$ :=$:=\set{\tau}{\tau\in\mathcal{D},r(\psi,\tau,t^{\prime})<0,\forall t^{\prime}\in t+I}.$ (15)
Let $\mathcal{T}_{p}=\{(\tau_{i},l_{i})|l_{i}=1\}_{i=1:N_{\mathrm{data}}}$ be the positive set of trajectories and $\mathcal{T}_{n}=T\setminus T_{p}$ be the negative set. For each predicate $\psi\in\Psi$, we then calculate its robustness distributions for $\mathcal{T}_{p}$ and $\mathcal{T}_{n}$$d(\mathcal{T}_{p},\psi)$ and $d(\mathcal{T}_{n},\psi)$, respectively—and their cosine similarity $S:=\frac{d(\mathcal{T}_{p},\psi)\cdot d(\mathcal{T}_{n},\psi)}{\lvert d(\mathcal{T}_{p},\psi)\rvert\lvert d(\mathcal{T}_{n},\psi)\rvert}$. As a result, predicates (and their negation) with a similarity score exceeding a user-defined threshold $S_{\mathrm{th}}$ are removed from $\Psi$ (and $\Psi_{\neg}$) prior to optimization.
[…]
$R_{\mathrm{T}}(\mathbf{w}^{\mathcal{F}},\mathbf{w}^{\mathcal{G}}):=\sum_{j=1}^{N_{\mathrm{AP}}}\max\left(\sum_{i=1}^{N_{\mathrm{AP}}}w_{ij}^{\mathcal{F}},\sum_{i=1}^{N_{\mathrm{AP}}}w_{i(j+N_{\mathrm{AP}})}^{\mathcal{F}}\right)\\ \max\left(\sum_{i=1}^{N_{\mathrm{AP}}}w_{ij}^{\mathcal{G}},\sum_{i=1}^{N_{\mathrm{AP}}}w_{i(j+N_{\mathrm{AP}})}^{\mathcal{G}}\right).$ (16)
The intuition behind the max operation is to choose the larger sum of weights of a predicate and those of its negation. Next, we define disjunctive clause regularizer $R_{\mathrm{D}}$. Given a temporal clause with importance weights $\mathbf{w}\in\{\mathbf{w}^{\mathcal{F}},\mathbf{w}^{\mathcal{G}}\}$, $R_{\mathrm{D}}$ is defined as,
$R_{\mathrm{D}}(\mathbf{w}):=\sum_{i=1}^{N_{\mathrm{AP}}-1}\sum_{j=i+1}^{N_{\mathrm{AP}}}\sum_{k=1}^{2N_{\mathrm{AP}}}w_{ik}w_{jk},$ (17)
where $w_{ij},w_{jk}\in\mathbf{w}$. The intuition behind the formulation is to differentiate the used predicates in a disjunctive clause in CNF from the predicates in the other disjunctive clauses within a temporal clauses.
Putting these regularizers together, the overall loss $\tilde{\mathcal{L}}$ is
~$\tilde{\mathcal{L}}(\mathcal{L},\mathbf{w}^{\mathcal{F}},\mathbf{w}^{\mathcal{G}})=\mathcal{L}+\lambda_{R_{\mathrm{T}}}R_{\mathrm{T}}(\mathbf{w}^{\mathcal{F}},\mathbf{w}^{\mathcal{G}})\\ +\lambda_{R_{\mathrm{D}}}[R_{\mathrm{D}}(\mathbf{w}^{\mathcal{F}})+R_{\mathrm{D}}(\mathbf{w}^{\mathcal{G}})],$ (18)
where regularizer coefficients $\lambda_{R_{\mathrm{T}}},\lambda_{R_{\mathrm{D}}}\in\mathbb{R}_{\geq 0}$ are small real values, and $\mathcal{L}$ is a desired loss function. In this work, we use the loss $\mathcal{L}$ proposed by [18],
$\mathcal{L}(\phi)=\sum_{j=1}^{N_{\mathrm{data}}}\exp(-\zeta l_{j}r(\phi,\tau_{j},t))$ (19)
The weight pruning stage further simplifies the currently inferred wSTL specification by reducing the number of active predicates. This step proceeds in two phases. First, all weights with a value of zero are pruned, ensuring that these predicates are excluded from subsequent optimization steps. Second, among the remaining nonzero weights, the $N_{\mathrm{w}}\in\mathbb{Z}_{>0}$ smallest weights—where $N_{\mathrm{w}}$ is a user-specified parameter—are removed to eliminate the least contributing parts of the specification. After pruning, the remaining nonzero weights are normalized such that their sum equals one.
III-C Trajectory Sampling
As described in Section II, the neural network is trained as a binary classifier to distinguish between positive and negative trajectories. Positive trajectories are obtained by executing the target policy in the environment of interest. Alternatively, they can be sampled from existing datasets, depending on the application setting. Negative trajectories are generated using a random-walk policy in this work, serving as contrasting samples for the classifier. However, the choice of policy used for generating negative trajectories is flexible and can be adapted based on the user’s objectives or the specific characteristics of the environment.
III-D Explanability Evaluation Metrics
Beyond classification performance, TLNet emphasizes the human-interpretability and alignment of inferred wSTL explanations with underlying policy behavior. While alignment is ultimately subjective and often requires qualitative assessment, we introduce three quantitative metrics—conciseness, consistency, and strictness—to characterize key aspects of explanation quality.
Let $\phi$ be an wSTL explanation we want to evaluate and $N$ be the number of temporal clauses that we expect to appear in the explanation. For example, our specifications expect $N=2$ temporal clauses. First, we define a conciseness metric as,
$\mathrm{Conciseness}:=\frac{1}{N}\sum_{n=1}^{N}\begin{cases}\frac{1}{m_{n}}&\text{if $m_{n}\neq 0$},\\ 0&\text{otherwise},\end{cases}$
where $m_{n}$ is the number of predicates present in $n$-th temporal clause of $\phi$. Conciseness ranges from 0 to 1, with higher scores implying a shorter explanation with at least one predicate inside each expected temporal clause.
We now define a consistency metric. Let $\Phi=\{\phi_{k}\}_{k=1:K}$ be a set of $K$ explanations for the same policy (e.g., from different training seeds), and $\phi_{kn}$ be the $n$-th temporal clause ($n\in[1,N]$) expected to appear within $\phi_{k}$. We define a set of unique temporal clause structures $\Xi^{n}$ for a set of the $n$-th temporal clauses $\Set{\phi_{kn}}_{k=1:K}$. Each unique structure $\xi^{n}\in\Xi^{n}$ has $c_{\xi^{n}}$ temporal clauses satisfying $\xi^{n}$. Then, consistency of $\Phi$ is defined as
$\mathrm{Consistency}:=\frac{1}{N}\sum_{n=1}^{N}\begin{cases}\max_{\xi^{n}\in\Xi^{n}}\frac{c_{\xi^{n}}}{\lvert\Xi^{n}\rvert}&\text{if $\Xi^{n}\neq\emptyset$},\\ 0&\text{otherwise}.\end{cases}$
Consistency ranges from 0 to 1, with higher scores indicating greater agreement in clause structures across explanations, reflecting structural stability in learned explanations.
Finally, we introduce a strictness metric. Let $\phi_{n}$ be the $n$-th temporal clause of an explanation $\phi$ and $P$ be the number of potentially present predicates in $\phi$ (e.g., $P=2N_{\mathrm{AP}}$ in TLNet). Denote the numbers of conjunctions and disjunctions in $\phi_{n}$ as $C_{n}$ and $D_{n}$, respectively. Then, strictness is defined as,
$\mathrm{Strictness}:=\frac{1}{N}\sum_{n=1}^{N}\begin{cases}\frac{1}{P-C_{n}+D_{n}}&\text{if $\phi_{n}$ is not empty},\\ 0&\text{otherwise.}\end{cases}$
Strictness ranges from 0 to 1, with higher scores implying stricter logical expressions by rewarding conjunctions and penalizing disjunctions.
III-E Weight Distribution in wSTL
A key assumption in our neural network architecture is that weights can be distributed across Boolean and temporal clauses. We now prove that such distribution is theoretically sound. Recall that an arbitrary Boolean formula $\phi$ can be expressed in CNF as,
$\phi=\bigwedge_{i=1}^{n}\bigvee_{j=1}^{m_{i}}(\neg)\psi_{j},$ (20)
where $n,m_{i}\in\mathbb{Z}_{>0}$ and $\psi_{ij}$ are predicates. The normalized weights of an arbitrary CNF wSTL formula can be syntactically distributed across Boolean operators as follows.
Theorem 1 (Syntactic Weight Distributivity of Boolean Operators).
A normalized wSTL formula $\phi$ in CNF can be represented as follows,
$\phi=\bigwedge_{i=1}^{n}w_{i}\bigvee_{j=1}^{m_{i}}w_{ij}(\neg)\psi_{ij}\equiv\bigwedge_{i=1}^{n}\bigvee_{j=1}^{m_{i}}w_{ij}^{\prime}(\neg)\psi_{ij},$ (21)
where $n,m_{i}\in\mathbb{Z}_{>0}$, $w_{i},w_{ij}\in\mathbb{R}_{>0}$, $w_{ij}^{\prime}:=w_{i}w_{ij}$, $\sum_{i=1}^{n}w_{i}=\sum_{j=1}^{m_{i}}w_{ij}=1$, and $\psi_{ij}$ are predicates. Furthermore, the weights $w_{ij}^{\prime}$ satisfy $\sum_{i=1}^{n}\sum_{j=1}^{m_{i}}w_{ij}^{\prime}=1$.
Proof.
[…]
$\wedge w_{2}(w_{21}(\neg)\psi_{21}\vee...\vee w_{2m_{2}}(\neg)\psi_{2m_{2}})$
... $\wedge w_{n}(w_{n1}(\neg)\psi_{n1}\vee...\vee w_{nm_{n}}(\neg)\psi_{nm_{n}})$
$=(w_{1}w_{11}(\neg)\psi_{11}\vee...\vee w_{1}w_{1m_{1}}(\neg)\psi_{1m_{1}})$
$\wedge(w_{2}w_{21}(\neg)\psi_{21}\vee...\vee w_{2}w_{2m_{2}}(\neg)\psi_{2m_{2}})$
... $\wedge(w_{n}w_{n1}(\neg)\psi_{n1}\vee...\vee w_{n}w_{nm_{n}}(\neg)\psi_{nm_{n}})$
$=\bigwedge_{i=1}^{n}\bigvee_{j=1}^{m_{i}}w_{i}w_{ij}(\neg)\psi_{j}=\{\text{RHS}\}.$
IV Results
We perform a series of empirical experiments to compare our method to relevant baselines for robotic tasks. We consider three baselines: {outline} \1 A greedy pruning strategy proposed by [21], which independently removes individual network weights that do not affect classification accuracy post-optimization. \1 Top-$k$ pruning strategies (top-3 and top-5) used by [23, 20], which only retain the $k$ largest network weights. All baselines were implemented using the same neural network architecture as TLNet to ensure a consistent basis for comparison. For each policy we aimed to explain, we collected 500 positive trajectories using the trained policy and 500 negative trajectories using a random-walk policy. Each dataset was split 80% and 20% for training and evaluation, respectively.
(a) CtF environment.
(b) Obstructed fetch push.
(c) Chased robot navigation.
Figure 4: Screenshots from our environments.
IV-A Test Environments
We consider the three environments shown in Figure 4, which include discrete and continuous state-action spaces, navigation and manipulation tasks, and varied numbers of predicates.
IV-A1 Capture-the-flag (CtF)
CtF is a discrete grid-world environment with adversarial dynamics where two agents compete to capture each other’s flags, based on [16]. If both agents occupy adjacent cells in a given territory, the defending agent eliminates the other with 75% probability. We considered blue agent policies optimized for five scenarios. The first four considered different red agent heuristic policies (capture, fight, patrol, roomba) with the same reward function (capture: +1, kill: +0.5, being captured: $-1$, timestep penalty: $-0.01$). The fifth scenario (capture 0) used a capture red agent with the reward function modified to remove the kill reward (kill: 0). The red agent policies are defined as follows: capture aggressively targets the flag; fight pursues the blue agent; patrol defends a boundary region; and roomba switches between capturing and defending. In all red agent policies, the red agent performs random actions 25% of the time. We defined eight atomic predicates based on euclidean distances among the blue agent (“ba”), blue flag (“bf”), blue territory (“bt”), red agent (“ra”), red flag (“rf”), and obstacle (“ob”) and binary defeat (“df”) status $\{1,-1\}$ of the red agent: $\psi_{\mathrm{ba,bt}}$, $\psi_{\mathrm{ba,ob}}$, $\psi_{\mathrm{ba,ra}}$, $\psi_{\mathrm{ba,rf}}$, $\psi_{\mathrm{ra,bf}}$, $\psi_{\mathrm{ra,bt}}$, $\psi_{\mathrm{ra,df}}$, $\psi_{\mathrm{ra,ob}}$.
IV-A2 Fetch Push
We also consider a continuous control task based on the Fetch mobile manipulator [24]. The robot must push a blue block (“b”) to a red target (“t”) using its gripper (“g”), without disturbing a black obstacle (“o”). The state and action spaces are continuous ($\mathcal{S}\subset\mathbb{R}^{55},a\in\mathbb{R}^{4}$). Eight predicates were defined based on the distances between the gripper, block, target, and obstacle, including drop-related (“d”) conditions: $\psi_{\mathrm{gb}}$, $\psi_{\mathrm{gt}}$, $\psi_{\mathrm{go}}$, $\psi_{\mathrm{bt}}$, $\psi_{\mathrm{bo}}$, $\psi_{\mathrm{ot}}$, $\psi_{\mathrm{bd}}$, $\psi_{\mathrm{od}}$.
IV-A3 Robot Navigation
We extended the CarGoal-v0 environment [25] to create a more complex adversarial navigation task. The ego robot (“e”) must reach a green goal (“g”) while avoiding a violet chaser (“c”) moving with a fixed speed ($v_{c}=7.5\times 10^{-4}$). Static hazards (“h”) and movable vase (“v”) are present but not penalized. The state and action dimensions are $s\in\mathbb{R}^{88}$ and $a\in\mathbb{R}^{2}$, respectively. Ten distance-based predicates were defined, involving euclidean distances between the robot, the chaser, the goal and the hazards: $\psi_{\mathrm{ec}}$, $\psi_{\mathrm{eg}}$, $\psi_{\mathrm{eh}}$, $\psi_{\mathrm{ev}}$, $\psi_{\mathrm{gc}}$, $\psi_{\mathrm{gh}}$, $\psi_{\mathrm{gv}}$, $\psi_{\mathrm{hc}}$, $\psi_{\mathrm{hv}}$, $\psi_{\mathrm{vc}}$.
To test robustness to noise, we included predicates irrelevant to task success (e.g., $\psi_{\mathrm{ra,ob}}$, $\psi_{\mathrm{ot}}$, $\psi_{\mathrm{hv}}$) and evaluated whether TLNet could eliminate them in its explanations.
IV-B Policy Optimization
TLNet is a general framework to generate explanations of robot policies, though, we focused on policies optimized by deep reinforcement learning algorithms in this work. We used Stable-Baselines3 [26] for training all policies. For the CtF and navigation environments, PPO [27] was used; for the Fetch push task, we used SAC [28] with Hindsight Experience Replay (HER) [29]. Hyperparameters were taken from prior work [16, 30, 25]. Blue agent policies in CtF achieved over 97% success, except the roomba case, which reached 77.5% success in 1,000 evaluation episodes. For each policy, we performed 10 independent optimization runs with random initialization and collected classification accuracy and explanability metrics.
TABLE I: Baseline Comparison of Representative Generated Explanations.
Scenarios TLNet (Ours) Greedy Top-3 Top-5
CtF Capture
$0.5\mathcal{F}[1.0\psi_{\mathrm{ba,rf}}]\wedge$
$0.5\mathcal{G}[0.30\psi_{\mathrm{ba,rf}}\vee 0.70\neg\psi_{\mathrm{ra,bf}}]$
$0.5\mathcal{F}[0.26\psi_{\mathrm{ba,rf}}\wedge(0.25\psi_{\mathrm{ba,rf}}\vee 0.33\neg\psi_{\mathrm{ra,bt}})\wedge$
$(0.09\neg\psi_{\mathrm{ba,bt}}\vee 0.07\neg\psi_{\mathrm{ra,bt}})]\wedge$
$0.5\mathcal{G}[0.36\psi_{\mathrm{ba,rf}}\vee 0.64\neg\psi_{\mathrm{ra,bf}}]$
$\mathcal{F}[0.73\neg\psi_{\mathrm{ra,bt}}$
$\wedge 0.27\neg\psi_{\mathrm{ba,bt}}]$
$\mathcal{F}[0.83\neg\psi_{\mathrm{ra,bt}}$
$\wedge 0.17\psi_{\mathrm{ba,rf}}]$
CtF Capture 0
$0.5\mathcal{F}[1.0\psi_{\mathrm{ba,rf}}]\wedge$
$0.5\mathcal{G}[0.31\psi_{\mathrm{ba,rf}}\vee 0.69\neg\psi_{\mathrm{ra,bf}}]$
$0.5\mathcal{F}[0.08\psi_{\mathrm{ba,rf}}\wedge(0.40\psi_{\mathrm{ba,rf}}\vee 0.31\neg\psi_{\mathrm{ra,bt}})\wedge$
$(0.07\neg\psi_{\mathrm{ba,bt}}\vee 0.14\neg\psi_{\mathrm{ra,bt}})]\wedge 0.5\mathcal{G}[(0.33\psi_{\mathrm{ba,rf}}$
$\vee 0.58\neg\psi_{\mathrm{ra,bf}})\wedge(0.05\psi_{\mathrm{ba,rf}}\vee 0.04\neg\psi_{\mathrm{ra,bt}})]$
$\mathcal{F}[0.67\neg\psi_{\mathrm{ra,bt}}$
$\wedge 0.33\neg\psi_{\mathrm{ba,bt}}]$
$\mathcal{F}[0.83\neg\psi_{\mathrm{ra,bt}}$
$\wedge 0.17\psi_{\mathrm{ba,rf}}]$
CtF Fight
$0.5\mathcal{F}[1.0\psi_{\mathrm{ba,rf}}]\wedge$
$0.5\mathcal{G}[0.33\psi_{\mathrm{ba,rf}}\vee 0.67\neg\psi_{\mathrm{ba,ra}}]$
$0.5\mathcal{F}[0.77\psi_{\mathrm{ba,rf}}\wedge(0.15\psi_{\mathrm{ba,rf}}\vee 0.08\psi_{\mathrm{ra,bf}})]\wedge$
$0.5\mathcal{G}[(0.15\neg\psi_{\mathrm{ba,bt}}\vee 0.11\neg\psi_{\mathrm{ra,df}})\wedge 0.06\psi_{\mathrm{ba,rf}}$
$\vee 0.04\neg\psi_{\mathrm{ra,df}})\wedge(0.06\psi_{\mathrm{ba,rf}}\vee 0.06\neg\psi_{\mathrm{ba,bt}}\vee$
$0.04\neg\psi_{\mathrm{ra,df}})\wedge(0.02\psi_{\mathrm{ba,ra}}\vee 0.19\psi_{\mathrm{ba,rf}}\vee$
$0.16\neg\psi_{\mathrm{ba,bt}}\vee 0.11\neg\psi_{\mathrm{ra,df}})]$
$\mathcal{F}[1.0\psi_{\mathrm{ba,rf}}]$ $\mathcal{F}[1.0\psi_{\mathrm{ba,rf}}]$
CtF Patrol
$0.5\mathcal{F}[1.0\psi_{\mathrm{ba,rf}}]\wedge$
$0.5\mathcal{G}[0.52\psi_{\mathrm{ba,rf}}\vee 0.48\neg\psi_{\mathrm{ra,bt}}]$
$0.5\mathcal{F}[0.35\psi_{\mathrm{ba,rf}}\wedge(0.09\psi_{\mathrm{ba,rf}}\vee 0.55\psi_{\mathrm{ra,df}})]\wedge$
$0.5\mathcal{G}[(0.29\psi_{\mathrm{ba,rf}}\vee 0.04\psi_{\mathrm{ra,bf}}\vee 0.07\neg\psi_{\mathrm{ra,df}})\wedge$
$(0.30\psi_{\mathrm{ba,rf}}\vee 0.19\psi_{\mathrm{ra,bf}}\vee 0.02\neg\psi_{\mathrm{ba,bt}})\wedge$
$0.04\psi_{\mathrm{ra,bf}}\wedge 0.05\neg\psi_{\mathrm{ra,bf}}]$
$\mathcal{F}[1.0\psi_{\mathrm{ra,df}}]$ $\mathcal{F}[1.0\psi_{\mathrm{ra,df}}]$
CtF Roomba
$0.5\mathcal{F}[1.0\psi_{\mathrm{ba,rf}}]\wedge$
$0.5\mathcal{G}[0.35\psi_{\mathrm{ba,rf}}\vee 0.65\neg\psi_{\mathrm{ra,bf}}]$
$0.5\mathcal{F}[1.0\psi_{\mathrm{ba,rf}}]\wedge 0.5\mathcal{G}[0.21\neg\psi_{\mathrm{ra,bf}}\wedge 0.47\psi_{\mathrm{ba,rf}}$
$\wedge(0.04\neg\psi_{\mathrm{ra,bf}}\vee 0.05\neg\psi_{\mathrm{ra,df}})\wedge(0.18\psi_{\mathrm{ba,rf}}\vee$
$0.03\psi_{\mathrm{ra,df}}\vee 0.04\neg\psi_{\mathrm{ra,bf}}\vee 0.02\neg\psi_{\mathrm{ra,bt}})]$
$\mathcal{F}[1.0\psi_{\mathrm{ba,bt}}]$
$\mathcal{F}[0.48\neg\psi_{\mathrm{ba,bt}}$
$\wedge 0.52\psi_{\mathrm{ba,rf}}]$
Fetch Push
$0.5\mathcal{F}[1.0\psi_{\mathrm{bt}}]\wedge$
$0.5\mathcal{G}[0.41\psi_{\mathrm{gb}}\vee 0.59\psi_{\mathrm{bt}}]$
$0.5\mathcal{F}[0.74\psi_{\mathrm{bt}}\wedge(0.10\psi_{\mathrm{bt}}\vee 0.16\psi_{\mathrm{od}})]\wedge$
$0.5\mathcal{G}[1.0\neg\psi_{\mathrm{bd}}]$
$\mathcal{G}[1.0\psi_{\mathrm{bt}}]$ $\mathcal{G}[1.0\psi_{\mathrm{bt}}]$
Robot Navi. $0.5\mathcal{F}[1.0\psi_{\mathrm{eg}}]\wedge 0.5\mathcal{G}[1.0\neg\psi_{\mathrm{ec}}]$ $\mathcal{F}[1.0\psi_{\mathrm{eg}}]$ $\mathcal{F}[1.0\psi_{\mathrm{eg}}]$ $\mathcal{F}[1.0\psi_{\mathrm{eg}}]$
TABLE II: Baseline Comparison of Conciseness, Consistency, Strictness, and Accuracy Metrics.
Conciseness Consistency Strictness Accuracy
Scenarios Ours Greedy Top-3 Top-5 Ours Greedy Top-3 Top-5 Ours Greedy Top-3 Top-5 Ours Greedy Top-3 Top-5
CtF Capture 0.41 0.21 0.20 0.22 0.33 0.08 0.08 0.13 0.33 0.08 0.08 0.13 0.95 0.95 0.95 0.94
CtF Capture 0 0.42 0.21 0.20 0.30 0.45 0.11 0.15 0.24 0.45 0.11 0.15 0.24 0.95 0.95 0.95 0.95
CtF Fight 0.39 0.18 0.25 0.25 0.68 0.09 0.50 0.50 0.68 0.09 0.50 0.50 0.94 0.92 0.94 0.93
CtF Patrol 0.38 0.18 0.25 0.28 0.63 0.05 0.50 0.53 0.63 0.05 0.50 0.53 0.95 0.93 0.94 0.94
CtF Roomba 0.39 0.12 0.21 0.16 0.27 0.06 0.10 0.09 0.27 0.06 0.10 0.09 0.87 0.89 0.88 0.88
Fetch Push 0.42 0.31 0.25 0.25 1.00 0.28 0.50 0.50 1.00 0.28 0.50 0.50 0.99 0.97 0.97 0.97
Robot Navi. 0.48 0.22 0.25 0.25 0.68 0.15 0.50 0.50 0.68 0.15 0.50 0.50 0.97 0.96 0.96 0.96
IV-C Experimental Results
Table I presents representative explanations for each scenario, comparing our method with three baseline. Across all scenarios, TLNet consistently inferred both task ($\mathcal{F}$) and constraint ($\mathcal{G}$) clauses, while the top-$k$ methods often captured only one of the two. This structural incompleteness stems from the fact that top-$k$ pruning lacks guarantees on preserving clause-level composition, despite yielding concise formulas with high classification accuracy.
Our method produced well-structured explanations across environments. For instance, task clauses included $\psi_{\mathrm{ba,rf}}$ (“blue agent captures the red flag”) in CtF, and $\psi_{\mathrm{bt}}$ (“block reaches the target”) and $\psi_{\mathrm{eg}}$ (“robot reaches the target” in Fetch and robot navigation. Constraint clauses correctly included safety conditions like $\neg\psi_{\mathrm{ra,bf}}$ (“the red agent does not capture the blue flag”) or $\neg\psi_{\mathrm{bd}}$ (“the block is not dropped”), with the exception of the navigation domain, where structural guarantees were again lacking in the baselines.
As shown in Table II, our method outperformed all baselines in conciseness, consistency, and strictness, with only marginal differences in classification accuracy. On average, TLNet achieved up to 1.9$\times$ higher conciseness, 2.6$\times$ higher consistency, and 2.7$\times$ higher strictness. The lower accuracy in the CtF Roomba scenario can be attributed to the suboptimality of the learned policy, which reduced the separability between positive and negative trajectories. These results support the qualitative findings above, validating the explainability of our method’s outputs.
Across CtF scenarios, TLNet consistently captured the task $\psi_{\mathrm{ba,rf}}$. The capture and capture 0 scenarios yielded identical constraint structures: $\mathcal{G}(0.30\,\psi_{\mathrm{ba,rf}}\vee 0.70\,\neg\psi_{\mathrm{ra,bf}})$ and $\mathcal{G}(0.31\,\psi_{\mathrm{ba,rf}}\vee 0.69\,\neg\psi_{\mathrm{ra,bf}})$, reflecting their similar optimized behaviors despite reward differences. In contrast, the fight scenario reflected the learned avoidance strategy through the constraint $\mathcal{G}(0.33\,\psi_{\mathrm{ba,rf}}\vee 0.67\,\neg\psi_{\mathrm{ba,ra}})$, emphasizing avoidance of battles.
The inferred explanation for Fetch Push was: “With importance 0.5, eventually the block reaches the target; with importance 0.5, globally, the grip contacts the block (0.41) or the block is on the target (0.59)”. While the task was correctly captured, the constraint clause was only partially valid due to the use of smoothed min/max activations. Since the grip reaches the block early in the episode, negative robustness signals were underemphasized. Additionally, the predicate $\neg\psi_{\mathrm{go}}$ (“grip does not touch the obstacle”) was not inferred, likely due to minimal violations in both positive and random-walk trajectories. This illustrates a limitation of the binary classification formulation and the trajectory sampling strategy. Future work could address this by using adversarial or policy-guided sampling to generate more informative negative examples.
For the navigation task, TLNet inferred a clear and aligned explanation: “With importance 0.5, eventually the robot reaches the goal; with importance 0.5, globally, it avoids the chaser.” The method successfully ignored irrelevant predicates (e.g., $\psi_{\mathrm{hc}}$, $\psi_{\mathrm{hv}}$), demonstrating robustness against unnecessary input predicates and alignment with the tasks and constraints.
TABLE III: Ablation Study of CtF Capture Scenario.
Metrics Full No reg. No pruning No filter
Accuracy 0.951 0.954 0.776 0.947
Conciseness 0.408 0.396 0.147 0.289
Consistency 0.325 0.233 0.2 0.175
Strictness 0.117 0.117 0.063 0.106
Table III presents ablation results on the CtF Capture scenario. The full pipeline achieved the highest conciseness, consistency, and strictness. A variant without regularization slightly outperformed in accuracy, which is expected, as regularization biases optimization toward simpler explanations. However, the accuracy trade-off was minimal ($<0.01$), while consistency improved by 39.4% and conciseness by 3.0%. Removing the predicate filter significantly reduced conciseness (27.2%) and consistency (46.1%), highlighting its importance. The most impactful component was pruning, particularly in reducing explanation size—nearly tripling conciseness. Overall, the ablation study confirms that each step of our simplification process plays a vital role in producing interpretable, structurally sound explanations without compromising performance.
V Conclusion
We introduced a neuro-symbolic framework, TLNet, for generating weighted Signal Temporal Logic (wSTL) explanations of robot policies, addressing limitations of trajectory classification-based methods. Our approach directly explains policies while refining explanations through predicate filtering, loss regularization, and weight pruning. To assess quality beyond classification metrics, we proposed human-interpretability metrics—conciseness, consistency, and strictness. Experiments in three simulated robotics environments showed superior explanation explanability with high accuracy, with an ablation study confirming our simplification process’s effectiveness. By bridging neural policy learning and formal logic-based explanations, our work advances transparency in robotics, with future directions including more complex wSTL structures, human-in-the-loop refinement, and real-world applications.

### v2

@@
 Abstract
-Neural network-based policies have demonstrated success in many robotic applications, but often lack human-explanability, which poses challenges in safety-critical deployments.
-To address this, we propose a neuro-symbolic explanation framework that generates a weighted signal temporal logic (wSTL) specification to describe a robot policy in a interpretable form.
-Existing methods typically produce explanations that are verbose and inconsistent, which hinders explainability, and loose, which do not give meaningful insights into the underlying policy.
+Learning-based policies have demonstrated success in many robotic applications, but often lack explainability, which poses challenges in safety-critical deployments.
+To address this, we propose a neuro-symbolic explanation framework that generates a weighted signal temporal logic (wSTL) specification which describes a robot policy in a human-interpretable form.
+Existing methods typically produce explanations that are verbose and inconsistent, which hinders explainability, and loose, which limits meaningful insights into the underlying policy.
 We address these issues by introducing a simplification process consisting of predicate filtering, regularization, and iterative pruning.
@@
 Our method is validated in three simulated robotic environments, where it outperforms baselines in generating concise, consistent, and strict wSTL explanations without sacrificing classification accuracy.
-This work bridges neural network policy learning with formal methods, contributing to safer and more transparent decision-making in robotics.
+This work bridges policy learning with explainability through formal methods, contributing to safer and more transparent decision-making in robotics.
 I Introduction
@@
 However, these neural network policies are inherently black-box in nature, which poses significant concerns for safety-critical robotic applications, such as autonomous vehicles, where explainability is paramount.
-To address these concerns, various approaches have been explored to enhance explainability of learned policies [1, 2, 3], including post-hoc analysis techniques and methods to learn policies that are inherently explainable.
+To address these concerns, various approaches have been explored to enhance explainability of learned policies [1, 2], including post-hoc analysis techniques [3, 4] and methods to learn policies that are inherently explainable [5, 6].
 In this paper, we focus on post-hoc analysis to enable explanation of any given robot policy.
 While existing approaches, like saliency maps [7] and feature attribution [8], provide valuable information, they often lack language-based explanations, which limits human-interpretability.
-Transformer-based language models have recently shown promising results for providing language-based explanations [6, 7, 8], but they do not support formal verification of the output explanation, which limits their applicability in domains requiring strong safety and correctness assurances [9].
+Autoregressive language models have recently shown promising results for providing language-based explanations [9, 10, 11], but they do not support formal verification of the output explanation, which limits their applicability in domains requiring strong safety and correctness assurances [12].
 Formal methods offer an alternative that emphasizes verifiability.
@@
 Our proposed network, TLNet, addresses these concerns.
-Figure 2: We propose a neuro-symbolic approach for generating an explanation of a given robot policy using wSTL.
-Our network architecture, TLNet, includes a simplification process that ensures resulting explanations balance classification accuracy with conciseness, consistency, and strictness.
 The presence of real-valued weights in wSTL has enabled the use of neural network approaches for inferring wSTL specifications that explain given trajectory data.
 These methods have shown promising results, where they show high classification performance for various applications, including anomaly detection in naval surveillance [21], fault diagnosis [22], and weather forecasting [23].
-However, existing work only focuses on classification accuracy, which does not fully capture the actual explanability of learned specifications.
+However, existing work only focuses on classification accuracy, which does not fully capture the actual explainability of learned specifications.
 For example, as shown in Figure 1, these methods often produce specifications that are unnecessarily verbose (i.e., they include many unnecessary predicates), inconsistent across seeds, and/or loose, such that they are satisfied but do not provide useful insights (e.g., “catch-all” specifications that include many conjunctions or overly simplified specifications that do not fully capture the task or constraints).
@@
 Our method targets a subclass of wSTL structures with arbitrary conjunctions and disjunctions of predicates in temporal clauses.
-TLNet includes a multi-step simplification process that reduces specification verbosity and improves consistency, while ensuring strictness is maintained, as shown in Figure 2.
+TLNet includes a multi-step simplification process that reduces specification verbosity and improves consistency, while ensuring strictness is maintained.
 This process includes a predicate filtering step, a regularized loss function, and iterative weight pruning.
-We also provide theoretical justification for our wSTL-based neural network architecture, which essentially reduce learning parameters by the distributivity of importance weight, and introduce novel metrics for quantifying explainability—conciseness, consistency, and strictness—that complement standard classification accuracy, addressing a significant gap in current evaluation practices.
+We also provide theoretical justification for our wSTL-based neural network architecture, which essentially reduces learning parameters by the distributivity of importance weight, and introduce novel metrics for quantifying explainability—conciseness, consistency, and strictness—that complement standard classification accuracy, addressing a significant gap in current evaluation practices.
 We demonstrate our method in three simulated robotics environments that vary in continuity, dimensionality, task structure, and policy complexity, and show that it generally outperforms baselines with respect to classification accuracy and our proposed explainability metrics.
-TLNet bridges the gap between neural network-based policy learning and providing formal explanations of learned policies, offering a step toward safer and more transparent robotic systems.
+TLNet bridges the gap between neural network-based policy learning and provides formal explanations of learned policies, offering a step toward safer and more transparent robotic systems.
 II Preliminaries
-Let a trajectory (or a signal) of $n$-dimensional states be a discrete-time series $\tau=s_{0},s_{1},..,s_{H}$, where $s_{t}\in S$ is a state at timestep $t\in\mathbb{Z}_{>0}$, $S\subseteq\mathbb{R}^{n}$ is the state space, and $H$ is the time horizon.
+Let a trajectory (or a signal) of $n$-dimensional states be a discrete-time series $\tau=s_{0},s_{1},\dots,s_{H}$, where $s_{t}\in S$ is a state at timestep $t\in\mathbb{Z}_{>0}$, $S\subseteq\mathbb{R}^{n}$ is the state space, and $H$ is the time horizon of $\tau$.
 We also define an interval $I$ as $I:=[a,b]=\{k\in\mathbb{Z}_{\geq 0}|a\leq k\leq b;a,b\in\mathbb{Z}_{\geq 0}\}$, $\lvert I\rvert$ as the cardinality of $I$, and $t+I$ as the interval $[t+a,t+b]$.
 II-A Weighted Signal Temporal Logic
-We consider wSTL syntax [17] defined recursively as $\phi:=\top\mid\psi\mid\neg\phi\mid\bigwedge_{i=1}^{N}{}^{w}\phi_{i}\mid\bigvee_{i=1}^{N}{}^{w}\phi_{i}\mid\mathcal{G}_{I}^{w}\phi\mid\mathcal{F}_{I}^{w}\phi$, where $\top$ is the logical True value, $\psi:=f(s)\geq c$ is an atomic predicate given a real-valued function $f:S\rightarrow\mathbb{R}$ over $S$ with constant $c\in\mathbb{R}$, and $\neg$, $\wedge$, and $\vee$ are Boolean negation, conjunction, and disjunction operators, respectively.
+We consider wSTL syntax [20] defined recursively as $\phi:=\top\mid\psi\mid\neg\phi\mid\bigwedge_{i=1}^{N}{}^{w}\phi_{i}\mid\bigvee_{i=1}^{N}{}^{w}\phi_{i}\mid\mathcal{G}_{I}^{w}\phi\mid\mathcal{F}_{I}^{w}\phi$, where $\top$ is the logical True value, $\psi:=f(s)\geq c$ is an atomic predicate given a function $f:S\rightarrow\mathbb{R}$ with constant $c\in\mathbb{R}$, and $\neg$, $\wedge$, and $\vee$ are Boolean negation, conjunction, and disjunction operators, respectively.
 For conjunctive and disjunctive operators, we define the weights as $w:=[w_{i}]_{i=1:N}\in\mathbb{R}^{N}_{>0}$, where $N$ is the number of subformulae in the operator.
@@
 We discuss the reasons for this requirement in Section III-E.
-Given this requirement, we define normalized wSTL syntax recursively as $\phi:=\top\mid\psi\mid\neg\phi\mid\bigwedge_{i=1}^{N}{}^{w}\phi_{i}\mid\bigvee_{i=1}^{N}{}^{w}\phi_{i}\mid\mathcal{G}_{I}^{w}\phi\mid\mathcal{F}_{I}^{w}\phi_{i},$ where all terminology carries over from wSTL, except we further require $\sum_{i=1}^{N}w_{i}=1$ and $\sum_{k\in I}w_{k}=1$, where $0\leq w_{i},w_{k}\leq 1$.
+Given this requirement, we define normalized wSTL syntax recursively as $\phi:=\top\mid\psi\mid\neg\phi\mid\bigwedge_{i=1}^{N}{}^{w}\phi_{i}\mid\bigvee_{i=1}^{N}{}^{w}\phi_{i}\mid\mathcal{G}_{I}^{w}\phi\mid\mathcal{F}_{I}^{w}\phi_{i},$ where all terminology carries over from wSTL, except we further require $\sum_{i=1}^{N}w_{i}=\sum_{k\in I}w_{k}=1$, where $w_{i},w_{k}\in[0,1]$.
 We allow the weights to be zero because the zero weights do not influence the overall robustness [21].
@@
 $r\left(\bigwedge_{i=1}^{N}{}^{w}\phi_{i},\tau,t\right)$ ::=:= $\otimes^{\wedge}(w,[r(\phi_{i},\tau,t)]_{i=1:N}),$ (3)
-$r\left(\bigvee_{i=1}^{N}{}^{w}\phi_{i},\tau,t\right)$ ::=:= $\oplus^{\vee}(w,[r(\phi_{i},\tau,t)]_{i=1:N}).$ (4)
+$r\left(\bigvee_{i=1}^{N}{}^{w}\phi_{i},\tau,t\right)$ ::=:= $\oplus^{\vee}(w,[r(\phi_{i},\tau,t)]_{i=1:N}),$ (4)
 $r\left(\mathcal{G}^{w}_{I}\phi,\tau,t\right)$ ::=:= $\otimes^{\mathcal{G}}_{I}(w,[r(\phi,\tau,t^{\prime})]_{t^{\prime}\in t+I}),$ (5)
@@
 For notational simplicity, we interchangeably use $r(\phi,\tau,t)$ and $r_{\tau t}^{\phi}$.
-Building from [18], we define the aggregation functions as follows,
+Building from [21], we define the aggregation functions with a scaling constant $\sigma>0$ as follows,
 $\otimes^{\wedge}(w,[r_{\tau t}^{\phi_{i}}]_{i=1:N})$ $:=\frac{\sum_{i=1}^{N}w_{i}r_{\tau t}^{\phi_{i}}\exp(\frac{-r_{\tau t}^{\phi_{i}}}{\sigma})}{\sum_{i=1}^{N}w_{i}\exp(\frac{-r_{\tau t}^{\phi_{i}}}{\sigma})},$ (7)
 $\oplus^{\vee}(w,[r_{\tau t}^{\phi_{i}}]_{i=1:N})$ $:=\frac{\sum_{i=1}^{N}w_{i}r_{\tau t}^{\phi_{i}}\exp(\frac{r_{\tau t}^{\phi_{i}}}{\sigma})}{\sum_{i=1}^{N}w_{i}\exp(\frac{r_{\tau t}^{\phi_{i}}}{\sigma})},$ (8)
-$\otimes^{\mathcal{G}}[w,[r_{\tau t^{\prime}}^{\phi}]_{t^{\prime}\in t+I})$ $:=\frac{\sum_{t^{\prime}\in t+I}w_{t^{\prime}}r_{\tau t^{\prime}}^{\phi}\exp(\frac{-r_{\tau t^{\prime}}^{\phi}}{\sigma})}{\sum_{t^{\prime}\in t+I}w_{t^{\prime}}\exp(\frac{-r_{\tau t^{\prime}}^{\phi}}{\sigma})},$ (9)
-$\oplus^{\mathcal{F}}[w,[r_{\tau t^{\prime}}^{\phi}]_{t^{\prime}\in t+I})$ $:=\frac{\sum_{t^{\prime}\in t+I}w_{t^{\prime}}r_{\tau t^{\prime}}^{\phi}\exp(\frac{r_{\tau t^{\prime}}^{\phi}}{\sigma})}{\sum_{t^{\prime}\in t+I}w_{t^{\prime}}\exp(\frac{r_{\tau t^{\prime}}^{\phi}}{\sigma})},$ (10)
-where $\sigma>0$ is a scaling constant.
+$\otimes^{\mathcal{G}}(w,[r_{\tau t^{\prime}}^{\phi}]_{t^{\prime}\in t+I})$ $:=\frac{\sum_{t^{\prime}\in t+I}w_{t^{\prime}}r_{\tau t^{\prime}}^{\phi}\exp(\frac{-r_{\tau t^{\prime}}^{\phi}}{\sigma})}{\sum_{t^{\prime}\in t+I}w_{t^{\prime}}\exp(\frac{-r_{\tau t^{\prime}}^{\phi}}{\sigma})},$ (9)
+$\oplus^{\mathcal{F}}(w,[r_{\tau t^{\prime}}^{\phi}]_{t^{\prime}\in t+I})$ $:=\frac{\sum_{t^{\prime}\in t+I}w_{t^{\prime}}r_{\tau t^{\prime}}^{\phi}\exp(\frac{r_{\tau t^{\prime}}^{\phi}}{\sigma})}{\sum_{t^{\prime}\in t+I}w_{t^{\prime}}\exp(\frac{r_{\tau t^{\prime}}^{\phi}}{\sigma})}.$ (10)
+We use $\sigma=0.5$ in this work.
+Figure 2: We propose a neuro-symbolic approach for generating an explanation of a given robot policy using wSTL.
+Our network architecture, TLNet, includes a simplification process that ensures resulting explanations balance classification accuracy with conciseness, consistency, and strictness.
 II-C Problem Statement
 We consider the following binary classification task.
-Let $\mathcal{T}=\Set{(\tau_{i},l_{i})}{i=1,\dots,N_{\mathrm{data}}}$ be a labeled dataset, where $\tau_{i}\in\mathbb{R}^{n\times H}$ is the $i^{th}$ trajectory with label $l_{i}\in\{1,-1\}$ and $N_{\mathrm{data}}\in\mathbb{Z}_{>0}$ is the number of data points in the dataset.
+Let $\mathcal{T}=\Set{(\tau_{i},l_{i})}_{i=1:N_{\mathrm{data}}}$ be a labeled dataset, where $\tau_{i}\in\mathbb{R}^{n\times H}$ is the $i^{th}$ trajectory with label $l_{i}\in\{1,-1\}$ and $N_{\mathrm{data}}\in\mathbb{Z}_{>0}$ is the number of data points in the dataset.
 Problem 1.
@@
 We first introduce our neural network architecture, TLNet, which allows us to represent wSTL structures in a fully differentiable form.
-We then introduce our simplification process, which enhances the explainability of the generated explanations, and our trajectory sampling process.
-Finally, we propose novel evaluation metrics for explanability—conciseness, consistency, and strictness—and provide theoretical justification for our network architecture.
-Algorithm 1 summarizes our overall approach.
+We then introduce our simplification process, which enhances the explainability of inferred explanations, and our trajectory sampling process.
+Finally, we propose novel evaluation metrics for explainability—conciseness, consistency, and strictness—and provide theoretical justification for our network architecture.
+Figure 2 and Algorithm 1 summarize our overall approach.
+III-A Neural Network Architecture
+Let $\Psi=\Set{\psi_{i}}_{i=1:N_{\mathrm{AP}}}$ be a predefined set of atomic predicates and $\Psi_{\neg}=\Set{\neg\psi_{i}}_{i=1:N_{\mathrm{AP}}}$ be a set of their negations.
+We focus on representing normalized wSTL structures composed of temporal operators containing these predicates in conjunctive normal form (CNF), as follows,
+ϕ=italic-$\phi=$ $0.5\mathcal{F}_{I}^{\bar{w}}\left[\bigwedge_{i=1}^{N_{\mathrm{AP}}}w_{i}^{\mathcal{F}}\bigvee_{j=1}^{2N_{\mathrm{AP}}}w_{ij}^{\mathcal{F}}\psi_{j}\right]$
+$\wedge$ $0.5\mathcal{G}_{I}^{\bar{w}}\left[\bigwedge_{i=1}^{N_{\mathrm{AP}}}w_{i}^{\mathcal{G}}\bigvee_{j=1}^{2N_{\mathrm{AP}}}w_{ij}^{\mathcal{G}}\psi_{j}\right],$ (11)
+where $w_{i}^{\mathcal{F}},w_{ij}^{\mathcal{F}},w_{i}^{\mathcal{G}},w_{ij}^{\mathcal{G}}\in[0,1]$ are importance weights and $\psi_{j}\in\Psi\cup\Psi_{\neg}$ are predefined atomic predicates.
+We define the weights $\bar{w}$ for temporal operators over interval $I=[0,H]$ to be $\bar{w}=[\bar{w}_{k}=1/\lvert I\rvert]_{k=1:\lvert I\rvert}$ for trajectory $\tau$.
+We define the conjunction weights of the temporal clauses to be 0.5 to equally weigh the importance of the task ($\mathcal{F}$) and constraint ($\mathcal{G}$) parts of the explanation.
+We choose this class of structures because they allow us to represent a general class of robot policies aiming to complete a task while satisfying a set of constraints [26, 19].
+We represent specifications of the form given in Equation 11 by designing a neural network $f_{\mathbf{w}}$ parameterized by $\mathbf{w}$, where $\mathbf{w}$ defines the importance weights of the currently represented specification $\phi_{\mathbf{w}}$.
+The input to the neural network is a sequence of the robustness values for each predicate $\psi\in\Psi\cup\Psi_{\neg}$ for each timestep in a trajectory $\tau$.
+The output is the robustness of the current specification, $r(\phi_{\mathbf{w}},\tau)$, for that trajectory.
+The neural network is trained to minimize a regularized classification loss (discussed in Section III-B), given a labeled dataset $\mathcal{T}$.
 Algorithm 1 Explanation Generation
-1:Dataset $\mathcal{T}$, predicates $\Psi=\Set{\psi_{i}}{i=1,\dots,N_{\mathrm{AP}}}$, similarity threshold $S_{\mathrm{th}}$, number of pruning iterations $N_{\mathrm{pr}}$, number of weights to prune $N_{\mathbf{w}}$, loss function $\tilde{\mathcal{L}}.$
+1:Dataset $\mathcal{T}$, predicates $\Psi=\Set{\psi_{i}}_{i=1:N_{\mathrm{AP}}}$, similarity threshold $S_{\mathrm{th}}$, number of pruning iterations $N_{\mathrm{pr}}$, number of weights to prune $N_{\mathbf{w}}$, loss function $\tilde{\mathcal{L}}.$
 2:Explanation $\phi$
@@
 10:     Optimize $\mathbf{w}$ to minimize $\tilde{\mathcal{L}}$.
-11:end for
+11:     if only one non-zero weight in $\mathbf{w}^{\mathcal{F}}$/$\mathbf{w}^{\mathcal{G}}$ then break.
 12:return $\phi$ generated from $\mathbf{w}$ and $\Psi$ using Equation 12.
-III-A Neural Network Architecture
-Let $\Psi=\Set{\psi_{i}}{i=1,\dots,N_{\mathrm{AP}}}$ be a predefined set of atomic predicates and $\Psi_{\neg}=\Set{\neg\psi_{i}}{i=1,\dots,N_{\mathrm{AP}}}$ be a set of their negations.
-We focus on representing wSTL structures composed of temporal operators containing these predicates in conjunctive normal form (CNF), as follows,
-$\phi=0.5\mathcal{F}_{I}^{w^{U}}\left[\bigwedge_{i=1}^{N_{\mathrm{AP}}}\bigvee_{j=1}^{2N_{\mathrm{AP}}}w_{ij}^{\mathcal{F}}\psi_{j}\right]\\ \wedge 0.5\mathcal{G}_{I}^{w^{U}}\left[\bigwedge_{i=1}^{N_{\mathrm{AP}}}\bigvee_{j=1}^{2N_{\mathrm{AP}}}w_{ij}^{\mathcal{G}}\psi_{j}\right],$ (11)
-where $w_{ij}^{\mathcal{F}},w_{ij}^{\mathcal{G}}\in[0,1]$ are importance weights and $\psi_{j}\in\Psi\cup\Psi_{\neg}$ are predefined atomic predicates defining specification $\phi$.
-We choose this class of structures because they allow us to represent a general class of robot policies aiming to complete a task while satisfying a set of constraints [22, 16].
-We define the weights $w^{U}$ for temporal operators over interval $I=[0,H]$ to be $w^{U}=[w^{U}_{k}=1/\lvert I\rvert]_{k=1:\lvert I\rvert}$ for trajectory $\tau$.
-We define the conjunction weights of the temporal clauses to be 0.5 to equally weigh the importance of the task ($\mathcal{F}$) and constraint ($\mathcal{G}$) parts of the explanation.
-The architecture of TLNet is designed to ensure the built-in architectural interpretability of the network weights, achieved by the bijective correspondence of the network weights and the importance weights of Equation 11.
-In Section III-E, we detail the theoretical justification of the formula manipulation to aggregate the weights of the conjunctive and disjunctive clauses in Equation 11.
-The output of TLNet is the robustness of the input trajectory with respect to the specification defined by the current network weights.
+We reduce the number of parameters that are optimized by distributing the weights in Equation 11, resulting in a specification of the following form,
+[…]
+$0.5\mathcal{F}_{I}^{\bar{w}}\left[\bigwedge_{i=1}^{N_{\mathrm{AP}}}\bigvee_{j=1}^{2N_{\mathrm{AP}}}\tilde{w}_{ij}^{\mathcal{F}}\psi_{j}\right]\wedge 0.5\mathcal{G}_{I}^{\bar{w}}\left[\bigwedge_{i=1}^{N_{\mathrm{AP}}}\bigvee_{j=1}^{2N_{\mathrm{AP}}}\tilde{w}_{ij}^{\mathcal{G}}\psi_{j}\right],$ (12)
+where $\mathbf{w}=\mathbf{w}^{\mathcal{F}}\mathbin{\|}\mathbf{w}^{\mathcal{G}}$, $\mathbf{w}^{\mathcal{F}}=[\tilde{w}_{ij}^{\mathcal{F}}]_{i=1:N_{\mathrm{AP}},j=1:2N_{\mathrm{AP}}}$, $\mathbf{w}^{\mathcal{G}}=[\tilde{w}_{ij}^{\mathcal{G}}]_{i=1:N_{\mathrm{AP}},j=1:2N_{\mathrm{AP}}}$, $\tilde{w}_{ij}^{\mathcal{F}}=w_{i}^{\mathcal{F}}w_{ij}^{\mathcal{F}}$, and $\tilde{w}_{ij}^{\mathcal{G}}=w_{i}^{\mathcal{F}}w_{ij}^{\mathcal{G}}$.
+We justify this weight distribution operation in Section III-E.
 An example TLNet architecture for a dataset containing two predicates is shown in Figure 3.
-The specification represented by this architecture is $\phi=0.5\mathcal{F}_{I}^{w^{U}}[(w_{11}^{\mathcal{F}}\psi_{1}\vee w_{12}^{\mathcal{F}}\psi_{2}\vee w_{13}^{\mathcal{F}}\neg\psi_{1}\vee w_{14}^{\mathcal{F}}\neg\psi_{2})\wedge(w_{21}^{\mathcal{F}}\psi_{1}\vee w_{22}^{\mathcal{F}}\psi_{2}\vee w_{23}^{\mathcal{F}}\neg\psi_{1}\vee w_{24}^{\mathcal{F}}\neg\psi_{2})]\wedge 0.5\mathcal{G}_{I}^{w^{U}}[(w_{11}^{\mathcal{G}}\psi_{1}\vee w_{12}^{\mathcal{G}}\psi_{2}\vee w_{13}^{\mathcal{G}}\neg\psi_{1}\vee w_{14}^{\mathcal{G}}\neg\psi_{2})\wedge(w_{21}^{\mathcal{G}}\psi_{1}\vee w_{22}^{\mathcal{G}}\psi_{2}\vee w_{23}^{\mathcal{G}}\neg\psi_{1}\vee w_{24}^{\mathcal{G}}\neg\psi_{2})]$.
 Figure 3: Example architecture of the proposed neural network when there are two predicates to consider ($\psi_{1}$ and $\psi_{2}$).
@@
 This process consists of three steps: a predicate filter, loss regularization, and weight pruning.
-The predicate filter eliminates predicates that exhibit similar behavior across both positive and negative trajectory sets, as they contribute little to discriminating between those sets and may introduce unnecessary complexity.
-Given a dataset of trajectories $\mathcal{D}$ and a predicate $\psi\in\Psi$, we define the robustness distribution as
+III-B1 Predicate Filter
+As a pre-processing step, we implement a predicate filter that eliminates predicates that exhibit similar behavior across both positive and negative trajectory sets, as they contribute little to discriminating between those sets and may introduce unnecessary complexity.
+Given an arbitrary dataset of trajectories $\mathcal{D}$ and an atomic predicate $\psi$, we define the robustness distribution for $\psi$ as,
 $d(\mathcal{D},\psi)=\frac{1}{|\mathcal{D}|}\left[\lvert\mathcal{D}_{1}\rvert,\lvert\mathcal{D}_{2}\rvert,\lvert\mathcal{D}_{3}\rvert\right],$ (13)
-where
-$\mathcal{D}_{1}$ :=$:=\set{\tau}{\tau\in\mathcal{D},r(\psi,\tau,t^{\prime})\geq 0,\forall t^{\prime}\in t+I},$ (13)
+$\mathcal{D}_{1}$ :=$:=\set{\tau}{r(\psi,\tau,t)\geq 0\quad\forall t\in[0,H],\tau\in\mathcal{D}},$ (14)
 $\mathcal{D}_{2}$ $:=\mathcal{D}\setminus(\mathcal{D}_{1}\cup\mathcal{D}_{3}),$ (15)
-$\mathcal{D}_{3}$ :=$:=\set{\tau}{\tau\in\mathcal{D},r(\psi,\tau,t^{\prime})<0,\forall t^{\prime}\in t+I}.$ (15)
-Let $\mathcal{T}_{p}=\{(\tau_{i},l_{i})|l_{i}=1\}_{i=1:N_{\mathrm{data}}}$ be the positive set of trajectories and $\mathcal{T}_{n}=T\setminus T_{p}$ be the negative set.
-For each predicate $\psi\in\Psi$, we then calculate its robustness distributions for $\mathcal{T}_{p}$ and $\mathcal{T}_{n}$$d(\mathcal{T}_{p},\psi)$ and $d(\mathcal{T}_{n},\psi)$, respectively—and their cosine similarity $S:=\frac{d(\mathcal{T}_{p},\psi)\cdot d(\mathcal{T}_{n},\psi)}{\lvert d(\mathcal{T}_{p},\psi)\rvert\lvert d(\mathcal{T}_{n},\psi)\rvert}$.
-As a result, predicates (and their negation) with a similarity score exceeding a user-defined threshold $S_{\mathrm{th}}$ are removed from $\Psi$ (and $\Psi_{\neg}$) prior to optimization.
-[…]
-$R_{\mathrm{T}}(\mathbf{w}^{\mathcal{F}},\mathbf{w}^{\mathcal{G}}):=\sum_{j=1}^{N_{\mathrm{AP}}}\max\left(\sum_{i=1}^{N_{\mathrm{AP}}}w_{ij}^{\mathcal{F}},\sum_{i=1}^{N_{\mathrm{AP}}}w_{i(j+N_{\mathrm{AP}})}^{\mathcal{F}}\right)\\ \max\left(\sum_{i=1}^{N_{\mathrm{AP}}}w_{ij}^{\mathcal{G}},\sum_{i=1}^{N_{\mathrm{AP}}}w_{i(j+N_{\mathrm{AP}})}^{\mathcal{G}}\right).$ (16)
-The intuition behind the max operation is to choose the larger sum of weights of a predicate and those of its negation.
-Next, we define disjunctive clause regularizer $R_{\mathrm{D}}$.
-Given a temporal clause with importance weights $\mathbf{w}\in\{\mathbf{w}^{\mathcal{F}},\mathbf{w}^{\mathcal{G}}\}$, $R_{\mathrm{D}}$ is defined as,
-$R_{\mathrm{D}}(\mathbf{w}):=\sum_{i=1}^{N_{\mathrm{AP}}-1}\sum_{j=i+1}^{N_{\mathrm{AP}}}\sum_{k=1}^{2N_{\mathrm{AP}}}w_{ik}w_{jk},$ (17)
-where $w_{ij},w_{jk}\in\mathbf{w}$.
-The intuition behind the formulation is to differentiate the used predicates in a disjunctive clause in CNF from the predicates in the other disjunctive clauses within a temporal clauses.
-Putting these regularizers together, the overall loss $\tilde{\mathcal{L}}$ is
-~$\tilde{\mathcal{L}}(\mathcal{L},\mathbf{w}^{\mathcal{F}},\mathbf{w}^{\mathcal{G}})=\mathcal{L}+\lambda_{R_{\mathrm{T}}}R_{\mathrm{T}}(\mathbf{w}^{\mathcal{F}},\mathbf{w}^{\mathcal{G}})\\ +\lambda_{R_{\mathrm{D}}}[R_{\mathrm{D}}(\mathbf{w}^{\mathcal{F}})+R_{\mathrm{D}}(\mathbf{w}^{\mathcal{G}})],$ (18)
-where regularizer coefficients $\lambda_{R_{\mathrm{T}}},\lambda_{R_{\mathrm{D}}}\in\mathbb{R}_{\geq 0}$ are small real values, and $\mathcal{L}$ is a desired loss function.
-In this work, we use the loss $\mathcal{L}$ proposed by [18],
-$\mathcal{L}(\phi)=\sum_{j=1}^{N_{\mathrm{data}}}\exp(-\zeta l_{j}r(\phi,\tau_{j},t))$ (19)
-The weight pruning stage further simplifies the currently inferred wSTL specification by reducing the number of active predicates.
-This step proceeds in two phases.
-First, all weights with a value of zero are pruned, ensuring that these predicates are excluded from subsequent optimization steps.
-Second, among the remaining nonzero weights, the $N_{\mathrm{w}}\in\mathbb{Z}_{>0}$ smallest weights—where $N_{\mathrm{w}}$ is a user-specified parameter—are removed to eliminate the least contributing parts of the specification.
-After pruning, the remaining nonzero weights are normalized such that their sum equals one.
+$\mathcal{D}_{3}$ :=$:=\set{\tau}{r(\psi,\tau,t)<0\quad\forall t\in[0,H],\tau\in\mathcal{D},}.$ (16)
+That is, $\mathcal{D}_{1}$ is the set of trajectories where $\psi$ is always satisfied, $\mathcal{D}_{2}$ is the set where $\psi$ is satisfied at least once, and $\mathcal{D}_{3}$ is the set where $\psi$ is never satisfied.
+Now, let $\mathcal{T}_{p}=\{(\tau,l)|l=1,(\tau,l)\in\mathcal{T}\}$ be the positive set of trajectories in $\mathcal{T}$ and $\mathcal{T}_{n}=\mathcal{T}\setminus\mathcal{T}_{p}$ be the negative set.
+For each predicate $\psi\in\Psi$, we then calculate $d(\mathcal{T}_{p},\psi)$ and $d(\mathcal{T}_{n},\psi)$, and then calculate the cosine similarity $S$ between those distributions.
+We then remove predicates from the set $\Psi$ (and their negation from $\Psi_{\neg}$) which have a cosine similarity $S>S_{\mathrm{th}}$, where $S_{\mathrm{th}}$ is a user-defined threshold.
+III-B2 Loss Regularization
+The regularization step uses two regularizers to improve the conciseness of inferred explanations. The first regularizer, a temporal clause regularizer, $R_{\mathrm{T}}$, is defined as,
+$R_{\mathrm{T}}(\mathbf{w}):=\sum_{j=1}^{N_{\mathrm{AP}}}\max\left(\sum_{i=1}^{N_{\mathrm{AP}}}\tilde{w}_{ij}^{\mathcal{F}},\sum_{i=1}^{N_{\mathrm{AP}}}\tilde{w}_{i(j+N_{\mathrm{AP}})}^{\mathcal{F}}\right)\\ \max\left(\sum_{i=1}^{N_{\mathrm{AP}}}\tilde{w}_{ij}^{\mathcal{G}},\sum_{i=1}^{N_{\mathrm{AP}}}\tilde{w}s_{i(j+N_{\mathrm{AP}})}^{\mathcal{G}}\right).$ (17)
+Intuitively, this regularizer encourages the task and constraint clauses to contain different atomic predicates.
+The second regularizer, a disjunctive clause regularizer, $R_{\mathrm{D}}$, is defined as,
+$R_{\mathrm{D}}(\mathbf{w}^{\mathcal{O}}):=\sum_{i=1}^{N_{\mathrm{AP}}-1}\sum_{j=i+1}^{N_{\mathrm{AP}}}\sum_{k=1}^{2N_{\mathrm{AP}}}\tilde{w}^{\mathcal{O}}_{ik}\tilde{w}^{\mathcal{O}}_{jk},$ (18)
+where $\mathbf{w}^{\mathcal{O}}\in\{\mathbf{w}^{\mathcal{F}},\mathbf{w}^{\mathcal{G}}\}$ defines the weights within one of our temporal operators and $\tilde{w}^{\mathcal{O}}_{ij},\tilde{w}^{\mathcal{O}}_{jk}\in\mathbf{w}^{\mathcal{O}}$.
+Intuitively, this regularizer encourages disjunctive clauses within a temporal clause to contain different predicates.
+Given $(\tau,l)\in\mathcal{T}$, our overall loss, $\tilde{\mathcal{L}}$, is then,
+~$\tilde{\mathcal{L}}(\tau,l,\mathbf{w})=\mathcal{L}(\tau,l,\mathbf{w})+$ $\lambda_{R_{\mathrm{T}}}R_{\mathrm{T}}(\mathbf{w})$
+$+$ $\lambda_{R_{\mathrm{D}}}[R_{\mathrm{D}}(\mathbf{w}^{\mathcal{F}})+R_{\mathrm{D}}(\mathbf{w}^{\mathcal{G}})],$ (19)
+where $\mathcal{L}$ is a classification loss and $\lambda_{R_{\mathrm{T}}},\lambda_{R_{\mathrm{D}}}\in\mathbb{R}_{\geq 0}$ are regularizer hyperparameters.
+We use $\lambda_{R_{\mathrm{T}}}=0.01$ and $\lambda_{R_{\mathrm{D}}}=0.1$ for this work.
+We use the classification loss from [21],
+$\mathcal{L}(\tau,l,\mathbf{w})=\exp(-\zeta lr(\phi_{\mathbf{w}},\tau)),$ (20)
+where $\zeta\in\mathbb{R}_{>0}$ is a hyperparameter ($\zeta=1$ for this work).
+III-B3 Weight Pruning
+The weight pruning step further simplifies the currently inferred explanation by removing unnecessary weights.
+First, all weights with a value of zero are removed from the network.
+Then, among the remaining weights, the $N_{\mathrm{w}}\in\mathbb{Z}_{>0}$ smallest weights—where $N_{\mathrm{w}}$ is a user-specified hyperparameter—are removed to eliminate the least contributing parts of the specification.
+The remaining weights are then normalized.
+The pruning process terminates if there is only one non-zero weight in $\mathbf{w}^{\mathcal{F}}$ or $\mathbf{w}^{\mathcal{G}}$ to ensure the task-constraint structure of the inferred explanation.
 III-C Trajectory Sampling
-As described in Section II, the neural network is trained as a binary classifier to distinguish between positive and negative trajectories.
-Positive trajectories are obtained by executing the target policy in the environment of interest.
-Alternatively, they can be sampled from existing datasets, depending on the application setting.
-Negative trajectories are generated using a random-walk policy in this work, serving as contrasting samples for the classifier.
-However, the choice of policy used for generating negative trajectories is flexible and can be adapted based on the user’s objectives or the specific characteristics of the environment.
-III-D Explanability Evaluation Metrics
-Beyond classification performance, TLNet emphasizes the human-interpretability and alignment of inferred wSTL explanations with underlying policy behavior.
-While alignment is ultimately subjective and often requires qualitative assessment, we introduce three quantitative metrics—conciseness, consistency, and strictness—to characterize key aspects of explanation quality.
-Let $\phi$ be an wSTL explanation we want to evaluate and $N$ be the number of temporal clauses that we expect to appear in the explanation.
-For example, our specifications expect $N=2$ temporal clauses.
-First, we define a conciseness metric as,
+We obtain positive trajectories in $\mathcal{T}$ by executing the target policy in the environment.
+These trajectories can also be sampled from existing datasets.
+We obtain negative trajectories in $\mathcal{T}$ by executing a random-walk policy in the environment.
+III-D Explainability Evaluation Metrics
+Beyond classification performance, TLNet also emphasizes the human-interpretability of inferred explanations.
+We quantify this emphasis by introducing three explainability metrics—conciseness, consistency, and strictness.
+Let $\phi$ be a wSTL explanation we want to evaluate and $N$ be the number of temporal clauses that we expect to appear in the explanation.
+For example, in this work, we expect the inferred explanations to have one $\mathcal{F}$ clause and one $\mathcal{G}$ clause (thus $N=2$) because we assume there should be a task component and a constraint component.
+Now, let $\phi_{n}$ be the $n$-th expected temporal clause in $\phi$, where $\phi_{n}=\emptyset$ if that clause does not exist in $\phi$.
+Let the (unweighted) STL counterpart of $\phi_{n}$ be $\hat{\phi}_{n}$, where $n\in\{1,\dots N\}$.
+For example, in our work, if $\phi=\mathcal{G}(0.1\psi_{1}\wedge 0.9\psi_{2})$, then $\phi_{1}=\emptyset$, $\hat{\phi}_{1}=\emptyset$, $\phi_{2}=\mathcal{G}(0.1\psi_{1}\wedge 0.9\psi_{2})$, and $\hat{\phi}_{2}=\mathcal{G}(\psi_{1}\wedge\psi_{2})$.
+III-D1 Conciseness
+We define the conciseness of $\phi$ as,
 $\mathrm{Conciseness}:=\frac{1}{N}\sum_{n=1}^{N}\begin{cases}\frac{1}{m_{n}}&\text{if $m_{n}\neq 0$},\\ 0&\text{otherwise},\end{cases}$
-where $m_{n}$ is the number of predicates present in $n$-th temporal clause of $\phi$.
+where $m_{n}$ is the number of predicates present in $n$-th expected temporal clause $\phi_{n}$.
 Conciseness ranges from 0 to 1, with higher scores implying a shorter explanation with at least one predicate inside each expected temporal clause.
-We now define a consistency metric.
-Let $\Phi=\{\phi_{k}\}_{k=1:K}$ be a set of $K$ explanations for the same policy (e.g., from different training seeds), and $\phi_{kn}$ be the $n$-th temporal clause ($n\in[1,N]$) expected to appear within $\phi_{k}$.
-We define a set of unique temporal clause structures $\Xi^{n}$ for a set of the $n$-th temporal clauses $\Set{\phi_{kn}}_{k=1:K}$.
-Each unique structure $\xi^{n}\in\Xi^{n}$ has $c_{\xi^{n}}$ temporal clauses satisfying $\xi^{n}$.
-Then, consistency of $\Phi$ is defined as
-$\mathrm{Consistency}:=\frac{1}{N}\sum_{n=1}^{N}\begin{cases}\max_{\xi^{n}\in\Xi^{n}}\frac{c_{\xi^{n}}}{\lvert\Xi^{n}\rvert}&\text{if $\Xi^{n}\neq\emptyset$},\\ 0&\text{otherwise}.\end{cases}$
+III-D2 Consistency
+Let $\Phi=[\phi^{k}]_{k=1:K}$ be a sequence of $K$ explanations for the same policy (e.g., from different training seeds).
+Then, let $\Phi_{n}=[\phi_{n}^{k}]_{k=1:K}$ be the sequence of the $n$-th expected temporal clauses in $\Phi$, where $\phi_{n}^{k}$ is the $n$-th expected temporal clause for $\phi^{k}$.
+Let the STL counterpart of $\Phi_{n}$ be $\hat{\Phi}_{n}$.
+Then, let the set of unique elements of $\hat{\Phi}_{n}$ be $\Sigma_{n}=\{\hat{\phi}_{n}\in\hat{\Phi}_{n}\}$.
+We define the consistency of $\Phi$ as,
+$\mathrm{Consistency}:=$
+$\frac{1}{N}\sum_{n=1}^{N}\begin{cases}\max_{\sigma_{n}\in\Sigma_{n}}\sum_{\hat{\phi}_{n}\in\hat{\Phi}_{n}}\frac{\mathbf{1}_{\sigma_{n}}(\hat{\phi}_{n})}{K\lvert\Sigma_{n}\rvert}&\text{if $\Sigma_{n}\neq\emptyset$},\\ 0&\text{otherwise},\end{cases}$
+where $\mathbf{1}_{\sigma_{n}}(\cdot)$ is the indicator function for $\sigma_{n}$.
 Consistency ranges from 0 to 1, with higher scores indicating greater agreement in clause structures across explanations, reflecting structural stability in learned explanations.
-Finally, we introduce a strictness metric.
-Let $\phi_{n}$ be the $n$-th temporal clause of an explanation $\phi$ and $P$ be the number of potentially present predicates in $\phi$ (e.g., $P=2N_{\mathrm{AP}}$ in TLNet).
+III-D3 Strictness
+Finally, let $P$ be the number of potentially present predicates in $\phi$ (e.g., $P=2N_{\mathrm{AP}}$ in TLNet).
 Denote the numbers of conjunctions and disjunctions in $\phi_{n}$ as $C_{n}$ and $D_{n}$, respectively.
-Then, strictness is defined as,
-$\mathrm{Strictness}:=\frac{1}{N}\sum_{n=1}^{N}\begin{cases}\frac{1}{P-C_{n}+D_{n}}&\text{if $\phi_{n}$ is not empty},\\ 0&\text{otherwise.}\end{cases}$
+We define the strictness of $\phi$ as,
+$\mathrm{Strictness}:=\frac{1}{N}\sum_{n=1}^{N}\begin{cases}\frac{1}{P-C_{n}+D_{n}}&\text{if $\phi_{n}\neq\emptyset$},\\ 0&\text{otherwise.}\end{cases}$
 Strictness ranges from 0 to 1, with higher scores implying stricter logical expressions by rewarding conjunctions and penalizing disjunctions.
 III-E Weight Distribution in wSTL
-A key assumption in our neural network architecture is that weights can be distributed across Boolean and temporal clauses.
-We now prove that such distribution is theoretically sound.
-Recall that an arbitrary Boolean formula $\phi$ can be expressed in CNF as,
-$\phi=\bigwedge_{i=1}^{n}\bigvee_{j=1}^{m_{i}}(\neg)\psi_{j},$ (20)
-where $n,m_{i}\in\mathbb{Z}_{>0}$ and $\psi_{ij}$ are predicates. The normalized weights of an arbitrary CNF wSTL formula can be syntactically distributed across Boolean operators as follows.
+A key assumption in TLNet is that the weights can be distributed across Boolean and temporal clauses.
+We now prove that such a distribution is theoretically sound.
+The normalized weights of a wSTL formula in CNF can be syntactically distributed across Boolean operators as follows.
 Theorem 1 (Syntactic Weight Distributivity of Boolean Operators).
 A normalized wSTL formula $\phi$ in CNF can be represented as follows,
-$\phi=\bigwedge_{i=1}^{n}w_{i}\bigvee_{j=1}^{m_{i}}w_{ij}(\neg)\psi_{ij}\equiv\bigwedge_{i=1}^{n}\bigvee_{j=1}^{m_{i}}w_{ij}^{\prime}(\neg)\psi_{ij},$ (21)
-where $n,m_{i}\in\mathbb{Z}_{>0}$, $w_{i},w_{ij}\in\mathbb{R}_{>0}$, $w_{ij}^{\prime}:=w_{i}w_{ij}$, $\sum_{i=1}^{n}w_{i}=\sum_{j=1}^{m_{i}}w_{ij}=1$, and $\psi_{ij}$ are predicates.
-Furthermore, the weights $w_{ij}^{\prime}$ satisfy $\sum_{i=1}^{n}\sum_{j=1}^{m_{i}}w_{ij}^{\prime}=1$.
+$\phi=\bigwedge_{i=1}^{n}w_{i}\bigvee_{j=1}^{m_{i}}w_{ij}(\neg)\psi_{ij}\equiv\bigwedge_{i=1}^{n}\bigvee_{j=1}^{m_{i}}\tilde{w}_{ij}(\neg)\psi_{ij},$ (21)
+where $n,m_{i}\in\mathbb{Z}_{>0}$, $w_{i},w_{ij}\in\mathbb{R}_{>0}$, $\tilde{w}_{ij}:=w_{i}w_{ij}$, $\sum_{i=1}^{n}w_{i}=\sum_{j=1}^{m_{i}}w_{ij}=1$, and $\psi_{ij}$ are predicates.
+Furthermore, the weights $\tilde{w}_{ij}$ satisfy $\sum_{i=1}^{n}\sum_{j=1}^{m_{i}}\tilde{w}_{ij}=1$.
 Proof.
 […]
-$\wedge w_{2}(w_{21}(\neg)\psi_{21}\vee...\vee w_{2m_{2}}(\neg)\psi_{2m_{2}})$
 ... $\wedge w_{n}(w_{n1}(\neg)\psi_{n1}\vee...\vee w_{nm_{n}}(\neg)\psi_{nm_{n}})$
 $=(w_{1}w_{11}(\neg)\psi_{11}\vee...\vee w_{1}w_{1m_{1}}(\neg)\psi_{1m_{1}})$
-$\wedge(w_{2}w_{21}(\neg)\psi_{21}\vee...\vee w_{2}w_{2m_{2}}(\neg)\psi_{2m_{2}})$
 ... $\wedge(w_{n}w_{n1}(\neg)\psi_{n1}\vee...\vee w_{n}w_{nm_{n}}(\neg)\psi_{nm_{n}})$
@@
 We perform a series of empirical experiments to compare our method to relevant baselines for robotic tasks.
-We consider three baselines: {outline} \1 A greedy pruning strategy proposed by [21], which independently removes individual network weights that do not affect classification accuracy post-optimization.
-\1 Top-$k$ pruning strategies (top-3 and top-5) used by [23, 20], which only retain the $k$ largest network weights.
+We consider three baselines for inference of wSTL explanations, each adapted to model our normalized wSTL structures (given by Equation 12): {outline} \1 NN-TLI [25]: this method implements a greedy pruning strategy that independently removes network weights that do not affect classification accuracy after optimization.
+We removed their weight-rounding process (to convert to STL) to consider the structure in Equation 12 and used the aggregation functions given in Figure 2, since our empirical results showed more training stability with those functions.
+\1 STONE (top-3) [21]: this method retains the 3 largest network weights after training.
+\1 STONE (top-5) [21, 24]: this method retains the 5 largest network weights after training.
 All baselines were implemented using the same neural network architecture as TLNet to ensure a consistent basis for comparison.
-For each policy we aimed to explain, we collected 500 positive trajectories using the trained policy and 500 negative trajectories using a random-walk policy.
+For each policy we explain, we collected 500 positive trajectories using the trained policy and 500 negative trajectories using a random-walk policy.
 Each dataset was split 80% and 20% for training and evaluation, respectively.
+Our code is available on GitHub after acceptance.
 (a) CtF environment.
 (b) Obstructed fetch push.
-(c) Chased robot navigation.
+(c) Robot Navigation.
 Figure 4: Screenshots from our environments.
 IV-A Test Environments
-We consider the three environments shown in Figure 4, which include discrete and continuous state-action spaces, navigation and manipulation tasks, and varied numbers of predicates.
+We consider the three environments shown in Figure 4, which include discrete and continuous state-action spaces, navigation and manipulation tasks, and varied predicates.
 IV-A1 Capture-the-flag (CtF)
@@
 We considered blue agent policies optimized for five scenarios.
-The first four considered different red agent heuristic policies (capture, fight, patrol, roomba) with the same reward function (capture: +1, kill: +0.5, being captured: $-1$, timestep penalty: $-0.01$).
+The first four considered different red agent heuristic policies (capture, fight, patrol, mixture) with the same reward function (capture: +1, kill: +0.5, being captured: $-1$, timestep penalty: $-0.01$).
 The fifth scenario (capture 0) used a capture red agent with the reward function modified to remove the kill reward (kill: 0).
-The red agent policies are defined as follows: capture aggressively targets the flag; fight pursues the blue agent; patrol defends a boundary region; and roomba switches between capturing and defending.
-In all red agent policies, the red agent performs random actions 25% of the time.
-We defined eight atomic predicates based on euclidean distances among the blue agent (“ba”), blue flag (“bf”), blue territory (“bt”), red agent (“ra”), red flag (“rf”), and obstacle (“ob”) and binary defeat (“df”) status $\{1,-1\}$ of the red agent: $\psi_{\mathrm{ba,bt}}$, $\psi_{\mathrm{ba,ob}}$, $\psi_{\mathrm{ba,ra}}$, $\psi_{\mathrm{ba,rf}}$, $\psi_{\mathrm{ra,bf}}$, $\psi_{\mathrm{ra,bt}}$, $\psi_{\mathrm{ra,df}}$, $\psi_{\mathrm{ra,ob}}$.
+The red agent policies are defined as follows: capture pursues the flag; fight pursues the blue agent; patrol defends a boundary region; and mixture switches between capturing and defending.
+Red agents perform a random action 25% of the time.
+We defined eight atomic predicates based on the defeated (“df”) status $\{1,-1\}$ of the red agent and euclidean distances among the blue agent (“ba”), blue flag (“bf”), blue territory (“bt”), red agent (“ra”), red flag (“rf”), and obstacle (“ob”): $\psi_{\mathrm{ba,bt}}$, $\psi_{\mathrm{ba,ob}}$, $\psi_{\mathrm{ba,ra}}$, $\psi_{\mathrm{ba,rf}}$, $\psi_{\mathrm{ra,bf}}$, $\psi_{\mathrm{ra,bt}}$, $\psi_{\mathrm{ra,df}}$, $\psi_{\mathrm{ra,ob}}$.
 IV-A2 Fetch Push
@@
 The robot must push a blue block (“b”) to a red target (“t”) using its gripper (“g”), without disturbing a black obstacle (“o”).
-The state and action spaces are continuous ($\mathcal{S}\subset\mathbb{R}^{55},a\in\mathbb{R}^{4}$).
+The state and action spaces are continuous ($s\in\mathbb{R}^{55},a\in\mathbb{R}^{4}$).
 Eight predicates were defined based on the distances between the gripper, block, target, and obstacle, including drop-related (“d”) conditions: $\psi_{\mathrm{gb}}$, $\psi_{\mathrm{gt}}$, $\psi_{\mathrm{go}}$, $\psi_{\mathrm{bt}}$, $\psi_{\mathrm{bo}}$, $\psi_{\mathrm{ot}}$, $\psi_{\mathrm{bd}}$, $\psi_{\mathrm{od}}$.
 IV-A3 Robot Navigation
-We extended the CarGoal-v0 environment [25] to create a more complex adversarial navigation task.
+Finally, we extended the CarGoal-v0 environment [28] to create a more complex adversarial navigation task.
 The ego robot (“e”) must reach a green goal (“g”) while avoiding a violet chaser (“c”) moving with a fixed speed ($v_{c}=7.5\times 10^{-4}$).
-Static hazards (“h”) and movable vase (“v”) are present but not penalized.
+Static hazards (“h”) and a movable vase (“v”) are present but not penalized.
 The state and action dimensions are $s\in\mathbb{R}^{88}$ and $a\in\mathbb{R}^{2}$, respectively.
-Ten distance-based predicates were defined, involving euclidean distances between the robot, the chaser, the goal and the hazards: $\psi_{\mathrm{ec}}$, $\psi_{\mathrm{eg}}$, $\psi_{\mathrm{eh}}$, $\psi_{\mathrm{ev}}$, $\psi_{\mathrm{gc}}$, $\psi_{\mathrm{gh}}$, $\psi_{\mathrm{gv}}$, $\psi_{\mathrm{hc}}$, $\psi_{\mathrm{hv}}$, $\psi_{\mathrm{vc}}$.
-To test robustness to noise, we included predicates irrelevant to task success (e.g., $\psi_{\mathrm{ra,ob}}$, $\psi_{\mathrm{ot}}$, $\psi_{\mathrm{hv}}$) and evaluated whether TLNet could eliminate them in its explanations.
+Ten distance-based predicates were defined, involving euclidean distances between the robot, the chaser, the goal, and the hazards: $\psi_{\mathrm{ec}}$, $\psi_{\mathrm{eg}}$, $\psi_{\mathrm{eh}}$, $\psi_{\mathrm{ev}}$, $\psi_{\mathrm{gc}}$, $\psi_{\mathrm{gh}}$, $\psi_{\mathrm{gv}}$, $\psi_{\mathrm{hc}}$, $\psi_{\mathrm{hv}}$, $\psi_{\mathrm{vc}}$.
+We included predicates irrelevant to task success in each environment (e.g., $\psi_{\mathrm{ra,ob}}$, $\psi_{\mathrm{ot}}$, $\psi_{\mathrm{hv}}$) to evaluate whether the considered methods could eliminate them from their explanations.
 IV-B Policy Optimization
@@
 Hyperparameters were taken from prior work [19, 33, 28].
-Blue agent policies in CtF achieved over 97% success, except the roomba case, which reached 77.5% success in 1,000 evaluation episodes.
-For each policy, we performed 10 independent optimization runs with random initialization and collected classification accuracy and explanability metrics.
-TABLE I: Baseline Comparison of Representative Generated Explanations.
-Scenarios TLNet (Ours) Greedy Top-3 Top-5
+For each policy, we performed 10 independent optimization runs with random initialization and collected classification accuracy and explainability metrics.
+The time horizon ($H$) for each environment is 100 for CtF, 200 for Fetch Push, and 400 for Robot Navigation.
+TABLE I: Comparison of the Most Frequently Inferred Explanations.
+Scenarios TLNet (Ours) NN-TLI STONE (top-3) STONE (top-5)
 CtF Capture
-$0.5\mathcal{F}[1.0\psi_{\mathrm{ba,rf}}]\wedge$
-$0.5\mathcal{G}[0.30\psi_{\mathrm{ba,rf}}\vee 0.70\neg\psi_{\mathrm{ra,bf}}]$
-$0.5\mathcal{F}[0.26\psi_{\mathrm{ba,rf}}\wedge(0.25\psi_{\mathrm{ba,rf}}\vee 0.33\neg\psi_{\mathrm{ra,bt}})\wedge$
+$0.5\mathcal{F}_{I}^{\bar{w}}[1.0\psi_{\mathrm{ba,rf}}]\wedge$
+$0.5\mathcal{G}_{I}^{\bar{w}}[0.30\psi_{\mathrm{ba,rf}}\vee 0.70\neg\psi_{\mathrm{ra,bf}}]$
+$0.5\mathcal{F}_{I}^{\bar{w}}[0.26\psi_{\mathrm{ba,rf}}\wedge(0.25\psi_{\mathrm{ba,rf}}\vee 0.33\neg\psi_{\mathrm{ra,bt}})\wedge$
 $(0.09\neg\psi_{\mathrm{ba,bt}}\vee 0.07\neg\psi_{\mathrm{ra,bt}})]\wedge$
-$0.5\mathcal{G}[0.36\psi_{\mathrm{ba,rf}}\vee 0.64\neg\psi_{\mathrm{ra,bf}}]$
-$\mathcal{F}[0.73\neg\psi_{\mathrm{ra,bt}}$
+$0.5\mathcal{G}_{I}^{\bar{w}}[0.36\psi_{\mathrm{ba,rf}}\vee 0.64\neg\psi_{\mathrm{ra,bf}}]$
+$\mathcal{F}_{I}^{\bar{w}}[0.73\neg\psi_{\mathrm{ra,bt}}$
 $\wedge 0.27\neg\psi_{\mathrm{ba,bt}}]$
-$\mathcal{F}[0.83\neg\psi_{\mathrm{ra,bt}}$
-$\wedge 0.17\psi_{\mathrm{ba,rf}}]$
-CtF Capture 0
-$0.5\mathcal{F}[1.0\psi_{\mathrm{ba,rf}}]\wedge$
-$0.5\mathcal{G}[0.31\psi_{\mathrm{ba,rf}}\vee 0.69\neg\psi_{\mathrm{ra,bf}}]$
-$0.5\mathcal{F}[0.08\psi_{\mathrm{ba,rf}}\wedge(0.40\psi_{\mathrm{ba,rf}}\vee 0.31\neg\psi_{\mathrm{ra,bt}})\wedge$
-$(0.07\neg\psi_{\mathrm{ba,bt}}\vee 0.14\neg\psi_{\mathrm{ra,bt}})]\wedge 0.5\mathcal{G}[(0.33\psi_{\mathrm{ba,rf}}$
-$\vee 0.58\neg\psi_{\mathrm{ra,bf}})\wedge(0.05\psi_{\mathrm{ba,rf}}\vee 0.04\neg\psi_{\mathrm{ra,bt}})]$
-$\mathcal{F}[0.67\neg\psi_{\mathrm{ra,bt}}$
-$\wedge 0.33\neg\psi_{\mathrm{ba,bt}}]$
-$\mathcal{F}[0.83\neg\psi_{\mathrm{ra,bt}}$
+$\mathcal{F}_{I}^{\bar{w}}[0.83\neg\psi_{\mathrm{ra,bt}}$
 $\wedge 0.17\psi_{\mathrm{ba,rf}}]$
 CtF Fight
-$0.5\mathcal{F}[1.0\psi_{\mathrm{ba,rf}}]\wedge$
-$0.5\mathcal{G}[0.33\psi_{\mathrm{ba,rf}}\vee 0.67\neg\psi_{\mathrm{ba,ra}}]$
-$0.5\mathcal{F}[0.77\psi_{\mathrm{ba,rf}}\wedge(0.15\psi_{\mathrm{ba,rf}}\vee 0.08\psi_{\mathrm{ra,bf}})]\wedge$
-$0.5\mathcal{G}[(0.15\neg\psi_{\mathrm{ba,bt}}\vee 0.11\neg\psi_{\mathrm{ra,df}})\wedge 0.06\psi_{\mathrm{ba,rf}}$
+$0.5\mathcal{F}_{I}^{\bar{w}}[1.0\psi_{\mathrm{ba,rf}}]\wedge$
+$0.5\mathcal{G}_{I}^{\bar{w}}[0.33\psi_{\mathrm{ba,rf}}\vee 0.67\neg\psi_{\mathrm{ba,ra}}]$
+$0.5\mathcal{F}_{I}^{\bar{w}}[0.77\psi_{\mathrm{ba,rf}}\wedge(0.15\psi_{\mathrm{ba,rf}}\vee 0.08\psi_{\mathrm{ra,bf}})]\wedge$
+$0.5\mathcal{G}_{I}^{\bar{w}}[(0.15\neg\psi_{\mathrm{ba,bt}}\vee 0.11\neg\psi_{\mathrm{ra,df}})\wedge 0.06\psi_{\mathrm{ba,rf}}$
 $\vee 0.04\neg\psi_{\mathrm{ra,df}})\wedge(0.06\psi_{\mathrm{ba,rf}}\vee 0.06\neg\psi_{\mathrm{ba,bt}}\vee$
@@
 $0.16\neg\psi_{\mathrm{ba,bt}}\vee 0.11\neg\psi_{\mathrm{ra,df}})]$
-$\mathcal{F}[1.0\psi_{\mathrm{ba,rf}}]$ $\mathcal{F}[1.0\psi_{\mathrm{ba,rf}}]$
+$\mathcal{F}_{I}^{\bar{w}}[1.0\psi_{\mathrm{ba,rf}}]$ $\mathcal{F}_{I}^{\bar{w}}[1.0\psi_{\mathrm{ba,rf}}]$
 CtF Patrol
-$0.5\mathcal{F}[1.0\psi_{\mathrm{ba,rf}}]\wedge$
-$0.5\mathcal{G}[0.52\psi_{\mathrm{ba,rf}}\vee 0.48\neg\psi_{\mathrm{ra,bt}}]$
-$0.5\mathcal{F}[0.35\psi_{\mathrm{ba,rf}}\wedge(0.09\psi_{\mathrm{ba,rf}}\vee 0.55\psi_{\mathrm{ra,df}})]\wedge$
-$0.5\mathcal{G}[(0.29\psi_{\mathrm{ba,rf}}\vee 0.04\psi_{\mathrm{ra,bf}}\vee 0.07\neg\psi_{\mathrm{ra,df}})\wedge$
+$0.5\mathcal{F}_{I}^{\bar{w}}[1.0\psi_{\mathrm{ba,rf}}]\wedge$
+$0.5\mathcal{G}_{I}^{\bar{w}}[0.52\psi_{\mathrm{ba,rf}}\vee 0.48\neg\psi_{\mathrm{ra,bt}}]$
+$0.5\mathcal{F}_{I}^{\bar{w}}[0.35\psi_{\mathrm{ba,rf}}\wedge(0.09\psi_{\mathrm{ba,rf}}\vee 0.55\psi_{\mathrm{ra,df}})]\wedge$
+$0.5\mathcal{G}_{I}^{\bar{w}}[(0.29\psi_{\mathrm{ba,rf}}\vee 0.04\psi_{\mathrm{ra,bf}}\vee 0.07\neg\psi_{\mathrm{ra,df}})\wedge$
 $(0.30\psi_{\mathrm{ba,rf}}\vee 0.19\psi_{\mathrm{ra,bf}}\vee 0.02\neg\psi_{\mathrm{ba,bt}})\wedge$
 $0.04\psi_{\mathrm{ra,bf}}\wedge 0.05\neg\psi_{\mathrm{ra,bf}}]$
-$\mathcal{F}[1.0\psi_{\mathrm{ra,df}}]$ $\mathcal{F}[1.0\psi_{\mathrm{ra,df}}]$
-CtF Roomba
-$0.5\mathcal{F}[1.0\psi_{\mathrm{ba,rf}}]\wedge$
-$0.5\mathcal{G}[0.35\psi_{\mathrm{ba,rf}}\vee 0.65\neg\psi_{\mathrm{ra,bf}}]$
-$0.5\mathcal{F}[1.0\psi_{\mathrm{ba,rf}}]\wedge 0.5\mathcal{G}[0.21\neg\psi_{\mathrm{ra,bf}}\wedge 0.47\psi_{\mathrm{ba,rf}}$
+$\mathcal{F}_{I}^{\bar{w}}[1.0\psi_{\mathrm{ra,df}}]$ $\mathcal{F}_{I}^{\bar{w}}[1.0\psi_{\mathrm{ra,df}}]$
+CtF Mixture
+$0.5\mathcal{F}_{I}^{\bar{w}}[1.0\psi_{\mathrm{ba,rf}}]\wedge$
+$0.5\mathcal{G}_{I}^{\bar{w}}[0.44\psi_{\mathrm{ba,rf}}\vee 0.56\neg\psi_{\mathrm{ra,bf}}]$
+$0.5\mathcal{F}_{I}^{\bar{w}}[1.0\psi_{\mathrm{ba,rf}}]\wedge 0.5\mathcal{G}_{I}^{\bar{w}}[0.21\neg\psi_{\mathrm{ra,bf}}\wedge 0.47\psi_{\mathrm{ba,rf}}$
 $\wedge(0.04\neg\psi_{\mathrm{ra,bf}}\vee 0.05\neg\psi_{\mathrm{ra,df}})\wedge(0.18\psi_{\mathrm{ba,rf}}\vee$
 $0.03\psi_{\mathrm{ra,df}}\vee 0.04\neg\psi_{\mathrm{ra,bf}}\vee 0.02\neg\psi_{\mathrm{ra,bt}})]$
-$\mathcal{F}[1.0\psi_{\mathrm{ba,bt}}]$
-$\mathcal{F}[0.48\neg\psi_{\mathrm{ba,bt}}$
+$\mathcal{F}_{I}^{\bar{w}}[1.0\psi_{\mathrm{ba,bt}}]$
+$\mathcal{F}_{I}^{\bar{w}}[0.48\neg\psi_{\mathrm{ba,bt}}$
 $\wedge 0.52\psi_{\mathrm{ba,rf}}]$
+CtF Capture 0
+$0.5\mathcal{F}_{I}^{\bar{w}}[1.0\psi_{\mathrm{ba,rf}}]\wedge$
+$0.5\mathcal{G}_{I}^{\bar{w}}[0.31\psi_{\mathrm{ba,rf}}\vee 0.69\neg\psi_{\mathrm{ra,bf}}]$
+$0.5\mathcal{F}_{I}^{\bar{w}}[0.08\psi_{\mathrm{ba,rf}}\wedge(0.40\psi_{\mathrm{ba,rf}}\vee 0.31\neg\psi_{\mathrm{ra,bt}})\wedge$
+$(0.07\neg\psi_{\mathrm{ba,bt}}\vee 0.14\neg\psi_{\mathrm{ra,bt}})]\wedge 0.5\mathcal{G}_{I}^{\bar{w}}[(0.33\psi_{\mathrm{ba,rf}}$
+$\vee 0.58\neg\psi_{\mathrm{ra,bf}})\wedge(0.05\psi_{\mathrm{ba,rf}}\vee 0.04\neg\psi_{\mathrm{ra,bt}})]$
+$\mathcal{F}_{I}^{\bar{w}}[0.67\neg\psi_{\mathrm{ra,bt}}$
+$\wedge 0.33\neg\psi_{\mathrm{ba,bt}}]$
+$\mathcal{F}_{I}^{\bar{w}}[0.83\neg\psi_{\mathrm{ra,bt}}$
+$\wedge 0.17\psi_{\mathrm{ba,rf}}]$
 Fetch Push
-$0.5\mathcal{F}[1.0\psi_{\mathrm{bt}}]\wedge$
-$0.5\mathcal{G}[0.41\psi_{\mathrm{gb}}\vee 0.59\psi_{\mathrm{bt}}]$
-$0.5\mathcal{F}[0.74\psi_{\mathrm{bt}}\wedge(0.10\psi_{\mathrm{bt}}\vee 0.16\psi_{\mathrm{od}})]\wedge$
-$0.5\mathcal{G}[1.0\neg\psi_{\mathrm{bd}}]$
-$\mathcal{G}[1.0\psi_{\mathrm{bt}}]$ $\mathcal{G}[1.0\psi_{\mathrm{bt}}]$
+$0.5\mathcal{F}_{I}^{\bar{w}}[1.0\psi_{\mathrm{bt}}]\wedge$
+$0.5\mathcal{G}_{I}^{\bar{w}}[0.41\psi_{\mathrm{gb}}\vee 0.59\psi_{\mathrm{bt}}]$
+$0.5\mathcal{F}_{I}^{\bar{w}}[0.74\psi_{\mathrm{bt}}\wedge(0.10\psi_{\mathrm{bt}}\vee 0.16\psi_{\mathrm{od}})]\wedge$
+$0.5\mathcal{G}_{I}^{\bar{w}}[1.0\neg\psi_{\mathrm{bd}}]$
+$\mathcal{G}_{I}^{\bar{w}}[1.0\psi_{\mathrm{bt}}]$ $\mathcal{G}_{I}^{\bar{w}}[1.0\psi_{\mathrm{bt}}]$
 Robot Navi.
-$0.5\mathcal{F}[1.0\psi_{\mathrm{eg}}]\wedge 0.5\mathcal{G}[1.0\neg\psi_{\mathrm{ec}}]$ $\mathcal{F}[1.0\psi_{\mathrm{eg}}]$ $\mathcal{F}[1.0\psi_{\mathrm{eg}}]$ $\mathcal{F}[1.0\psi_{\mathrm{eg}}]$
-TABLE II: Baseline Comparison of Conciseness, Consistency, Strictness, and Accuracy Metrics.
+$0.5\mathcal{F}_{I}^{\bar{w}}[1.0\psi_{\mathrm{eg}}]\wedge 0.5\mathcal{G}_{I}^{\bar{w}}[1.0\neg\psi_{\mathrm{ec}}]$ $\mathcal{F}_{I}^{\bar{w}}[1.0\psi_{\mathrm{eg}}]$ $\mathcal{F}_{I}^{\bar{w}}[1.0\psi_{\mathrm{eg}}]$ $\mathcal{F}_{I}^{\bar{w}}[1.0\psi_{\mathrm{eg}}]$
+TABLE II: Comparison of Conciseness, Consistency, Strictness, and Accuracy Metrics. The best results are bolded.
 Conciseness Consistency Strictness Accuracy
-Scenarios Ours Greedy Top-3 Top-5 Ours Greedy Top-3 Top-5 Ours Greedy Top-3 Top-5 Ours Greedy Top-3 Top-5
+STONE STONE STONE STONE
+Scenarios Ours NN-TLI Top-3 Top-5 Ours NN-TLI Top-3 Top-5 Ours NN-TLI Top-3 Top-5 Ours NN-TLI Top-3 Top-5
 CtF Capture 0.41 0.21 0.20 0.22 0.33 0.08 0.08 0.13 0.33 0.08 0.08 0.13 0.95 0.95 0.95 0.94
-CtF Capture 0 0.42 0.21 0.20 0.30 0.45 0.11 0.15 0.24 0.45 0.11 0.15 0.24 0.95 0.95 0.95 0.95
 CtF Fight 0.39 0.18 0.25 0.25 0.68 0.09 0.50 0.50 0.68 0.09 0.50 0.50 0.94 0.92 0.94 0.93
 CtF Patrol 0.38 0.18 0.25 0.28 0.63 0.05 0.50 0.53 0.63 0.05 0.50 0.53 0.95 0.93 0.94 0.94
-CtF Roomba 0.39 0.12 0.21 0.16 0.27 0.06 0.10 0.09 0.27 0.06 0.10 0.09 0.87 0.89 0.88 0.88
+CtF Mixture 0.39 0.12 0.21 0.16 0.27 0.06 0.10 0.09 0.27 0.06 0.10 0.09 0.87 0.89 0.88 0.88
+CtF Capt. 0 0.42 0.21 0.20 0.30 0.45 0.11 0.15 0.24 0.45 0.11 0.15 0.24 0.95 0.95 0.95 0.95
 Fetch Push 0.42 0.31 0.25 0.25 1.00 0.28 0.50 0.50 1.00 0.28 0.50 0.50 0.99 0.97 0.97 0.97
@@
 IV-C Experimental Results
-Table I presents representative explanations for each scenario, comparing our method with three baseline.
-Across all scenarios, TLNet consistently inferred both task ($\mathcal{F}$) and constraint ($\mathcal{G}$) clauses, while the top-$k$ methods often captured only one of the two.
-This structural incompleteness stems from the fact that top-$k$ pruning lacks guarantees on preserving clause-level composition, despite yielding concise formulas with high classification accuracy.
-Our method produced well-structured explanations across environments.
-For instance, task clauses included $\psi_{\mathrm{ba,rf}}$ (“blue agent captures the red flag”) in CtF, and $\psi_{\mathrm{bt}}$ (“block reaches the target”) and $\psi_{\mathrm{eg}}$ (“robot reaches the target” in Fetch and robot navigation.
-Constraint clauses correctly included safety conditions like $\neg\psi_{\mathrm{ra,bf}}$ (“the red agent does not capture the blue flag”) or $\neg\psi_{\mathrm{bd}}$ (“the block is not dropped”), with the exception of the navigation domain, where structural guarantees were again lacking in the baselines.
-As shown in Table II, our method outperformed all baselines in conciseness, consistency, and strictness, with only marginal differences in classification accuracy.
-On average, TLNet achieved up to 1.9$\times$ higher conciseness, 2.6$\times$ higher consistency, and 2.7$\times$ higher strictness.
-The lower accuracy in the CtF Roomba scenario can be attributed to the suboptimality of the learned policy, which reduced the separability between positive and negative trajectories.
-These results support the qualitative findings above, validating the explainability of our method’s outputs.
-Across CtF scenarios, TLNet consistently captured the task $\psi_{\mathrm{ba,rf}}$.
-The capture and capture 0 scenarios yielded identical constraint structures: $\mathcal{G}(0.30\,\psi_{\mathrm{ba,rf}}\vee 0.70\,\neg\psi_{\mathrm{ra,bf}})$ and $\mathcal{G}(0.31\,\psi_{\mathrm{ba,rf}}\vee 0.69\,\neg\psi_{\mathrm{ra,bf}})$, reflecting their similar optimized behaviors despite reward differences.
-In contrast, the fight scenario reflected the learned avoidance strategy through the constraint $\mathcal{G}(0.33\,\psi_{\mathrm{ba,rf}}\vee 0.67\,\neg\psi_{\mathrm{ba,ra}})$, emphasizing avoidance of battles.
-The inferred explanation for Fetch Push was: “With importance 0.5, eventually the block reaches the target; with importance 0.5, globally, the grip contacts the block (0.41) or the block is on the target (0.59)”.
-While the task was correctly captured, the constraint clause was only partially valid due to the use of smoothed min/max activations.
-Since the grip reaches the block early in the episode, negative robustness signals were underemphasized.
-Additionally, the predicate $\neg\psi_{\mathrm{go}}$ (“grip does not touch the obstacle”) was not inferred, likely due to minimal violations in both positive and random-walk trajectories.
-This illustrates a limitation of the binary classification formulation and the trajectory sampling strategy.
-Future work could address this by using adversarial or policy-guided sampling to generate more informative negative examples.
-For the navigation task, TLNet inferred a clear and aligned explanation: “With importance 0.5, eventually the robot reaches the goal; with importance 0.5, globally, it avoids the chaser.” The method successfully ignored irrelevant predicates (e.g., $\psi_{\mathrm{hc}}$, $\psi_{\mathrm{hv}}$), demonstrating robustness against unnecessary input predicates and alignment with the tasks and constraints.
-TABLE III: Ablation Study of CtF Capture Scenario.
-Metrics Full No reg. No pruning No filter
-Accuracy 0.951 0.954 0.776 0.947
-Conciseness 0.408 0.396 0.147 0.289
-Consistency 0.325 0.233 0.2 0.175
-Strictness 0.117 0.117 0.063 0.106
-Table III presents ablation results on the CtF Capture scenario.
-The full pipeline achieved the highest conciseness, consistency, and strictness.
-A variant without regularization slightly outperformed in accuracy, which is expected, as regularization biases optimization toward simpler explanations.
-However, the accuracy trade-off was minimal ($<0.01$), while consistency improved by 39.4% and conciseness by 3.0%.
-Removing the predicate filter significantly reduced conciseness (27.2%) and consistency (46.1%), highlighting its importance.
-The most impactful component was pruning, particularly in reducing explanation size—nearly tripling conciseness.
-Overall, the ablation study confirms that each step of our simplification process plays a vital role in producing interpretable, structurally sound explanations without compromising performance.
+IV-C1 Benefits and Limitations of TLNet
+Table I presents the most frequent explanation obtained by each method for each scenario.
+Across all scenarios, TLNet consistently inferred both task and constraint clauses, while also producing concise and relatively easy-to-understand explanations.
+In comparison, NN-TLI typically yielded verbose explanations that are difficult to understand, while STONE often captured only one temporal clause, resulting in incomplete explanations.
+The explanations produced by TLNet are also qualitatively reasonable (and often correct).
+For example, for all CtF scenarios, TLNet correctly inferred the task clause as $\psi_{\mathrm{ba,rf}}$ (“blue agent captures the red flag”)—no other method correctly inferred this task for every scenario.
+The inferred CtF constraints are also syntactically reasonable.
+For example, in the Capture scenario, the blue agent (global) constraint includes “red agent not on the blue flag” because the capture red policy only targets the blue flag, which required the blue agent to learn to prevent the red agent from reaching the blue flag.
+Other scenarios, such as Fight and Patrol, did not include this predicate in the inferred constraint because those red policies do not target the blue flag.
+In comparison, the CtF constraints inferred by NN-TLI are generally too verbose to understand, while STONE did not infer any constraints.
+For Fetch Push, TLNet also correctly inferred the task clause as $\psi_{\mathrm{bt}}$ (“block reaches the target”), which no other method correctly inferred.
+While TLNet inferred part of the expected constraint ($\psi_{\mathrm{gb}},\psi_{\mathrm{bt}}$), it missed an expected constraint $\neg\psi_{\mathrm{go}}$ (“grip does not touch the obstacle”), likely due to insufficient occurrences of the grip touching the obstacle in the sampled data.
+Future research could mitigate this issue by introducing more intelligent sampling strategies for generating the negative dataset (e.g., adversarial or targeted trajectory sampling).
+NN-TLI also inferred part of the expected constraint, but also missed the $\neg\psi_{\mathrm{go}}$ constraint.
+Meanwhile, both variations of STONE inferred an incorrect constraint of $\psi_{\mathrm{bt}}$ (“block reaches the target”), since this is actually the task and not possible to always satisfy.
+For Robot Navigation, all methods correctly inferred the task clause but TLNet was the only one to infer the correct constraint clause as $\neg\psi_{\mathrm{ec}}$ (“the robot does not hit the chaser”).
+As shown in Table II, TLNet also outperformed all baselines in conciseness, consistency, and strictness, with marginal differences in classification accuracy.
+Across all scenarios, TLNet achieved up to 1.9$\times$ higher conciseness, 2.6$\times$ higher consistency, and 2.7$\times$ higher strictness.
+These results support the qualitative findings above, validating the explainability of our method’s outputs and the use of these explainability metrics.
+IV-C2 Benefit of Including Weights
+We now demonstrate the benefit of including weights within inferred explanations (e.g., using wSTL rather than STL [25] or LTL [16, 17, 18]) by examining how inferred weights differ across CtF scenarios.
+Consider the Capture and Mixture explanations in Table I.
+Both capture and mixture red policies target the blue flag, but the capture policy is more aggressive in doing so.
+The blue agent, in response, learns slightly different policies: a more defensive policy in Capture, prioritizing the protection of its flag, and a more aggressive policy in Mixture, focusing more on capturing the red flag.
+This policy difference is reflected in the weights in the constraint clause, where the $\neg\psi_{\mathrm{ra,bf}}$ predicate receives a higher weight in the Capture explanation (0.70) than in the Mixture explanation (0.56), due to the Capture policy’s prioritization of flag protection.
+Conversely, the Mixture policy shows a higher weight for the $\psi_{\mathrm{ba,rf}}$ predicate (0.44) than the Capture policy (0.30), due to the Mixture policy’s prioritization of flag capture.
+Note that $\psi_{\mathrm{ba,rf}}$ appears in both the task and constraint clauses in these explanations.
+This duplication results from our design choice to fix equal weights (0.5) on the task and constraint clauses, which forces TLNet to express prioritization between the task and constraint by including relevant task predicates in the constraint clause.
+Although one could alternatively include the task and constraints weights within the optimized parameters, we found that doing so introduced training instability.
+Now, consider the Capture and Capture 0 scenarios, which differ in the reward to kill the red agent (0.5 vs.
+0).
+While one might naively expect this difference to produce different policy explanations, it actually should not change the optimal blue agent policies because, to win the game, the blue agent should first defend its flag (by killing the red agent), since the red agent uses the capture policy.
+We see that TLNet successfully captured this nuance by inferring constraints with nearly identical weights, reflecting its robustness to minor reward differences that still lead to similar policies.
+NN-TLI and STONE did not infer similar constraints across these scenarios.
+More broadly, this result also demonstrates the importance of inferring explanations from actual observed behaviors of RL agents rather than simply looking at their specified reward function.
+TABLE III: Ablation Study of TLNet in the CtF Capture Scenario.
+Metrics Full No Filter No Regularization No Pruning
+Conciseness 0.408 0.289 0.396 0.147
+Consistency 0.325 0.175 0.233 0.2
+Strictness 0.117 0.106 0.117 0.063
+Accuracy 0.951 0.947 0.954 0.776
+IV-C3 Importance of TLNet Components
+Finally, the ablation study results shown in Table III show that filtering, regularization, and pruning are all important to the performance of TLNet.
+Removing predicate filtering notably reduced conciseness (by 27.2%) and consistency (by 46.1%).
+Removing regularization decreased conciseness (by 2.94%) and significanbly reduced consistency (by 28.3%) with marginal accuracy improvement.
+Finally, removing the pruning step nearly tripled conciseness and halved strictness.
 V Conclusion
-We introduced a neuro-symbolic framework, TLNet, for generating weighted Signal Temporal Logic (wSTL) explanations of robot policies, addressing limitations of trajectory classification-based methods.
-Our approach directly explains policies while refining explanations through predicate filtering, loss regularization, and weight pruning.
-To assess quality beyond classification metrics, we proposed human-interpretability metrics—conciseness, consistency, and strictness.
-Experiments in three simulated robotics environments showed superior explanation explanability with high accuracy, with an ablation study confirming our simplification process’s effectiveness.
-By bridging neural policy learning and formal logic-based explanations, our work advances transparency in robotics, with future directions including more complex wSTL structures, human-in-the-loop refinement, and real-world applications.
+We introduce a neuro-symbolic framework, based on the TLNet architecture, for generating wSTL explanations of robot policies.
+Our approach generates easy-to-understand explanations through predicate filtering, loss regularization, and weight pruning.
+To assess explanation quality beyond classification metrics, we also propose explainability metrics—conciseness, consistency, and strictness.
+Experiments in three simulated robotic environments show that our method produces explanations that better capture these metrics while retaining high classification accuracy.
+By bridging neural network learning and formal logic-based explanations, our work advances explainability in complex robotic applications, with future directions including more expressive wSTL structures, human-in-the-loop refinement, and explainability-by-design policy optimization.

### v3

@@
 Abstract
-Learning-based policies have demonstrated success in many robotic applications, but often lack explainability, which poses challenges in safety-critical deployments.
-To address this, we propose a neuro-symbolic explanation framework that generates a weighted signal temporal logic (wSTL) specification which describes a robot policy in a human-interpretable form.
-Existing methods typically produce explanations that are verbose and inconsistent, which hinders explainability, and loose, which limits meaningful insights into the underlying policy.
+Learning-based policies have demonstrated success in many robotic applications, but often lack explainability.
+We propose a neuro-symbolic explanation framework that generates a weighted signal temporal logic (wSTL) specification which describes a robot policy in a human-interpretable form.
+Existing methods typically produce explanations that are verbose and inconsistent, which hinders explainability, and are loose, which limits meaningful insights.
 We address these issues by introducing a simplification process consisting of predicate filtering, regularization, and iterative pruning.
-We also introduce three novel explainability evaluation metrics—conciseness, consistency, and strictness—to assess explanation quality beyond conventional classification metrics.
-Our method is validated in three simulated robotic environments, where it outperforms baselines in generating concise, consistent, and strict wSTL explanations without sacrificing classification accuracy.
-This work bridges policy learning with explainability through formal methods, contributing to safer and more transparent decision-making in robotics.
+We also introduce three explainability metrics—conciseness, consistency, and strictness—to assess explanation quality beyond conventional classification accuracy.
+Our method—TLNet—is validated in three simulated robotic environments, where it outperforms baselines in generating concise, consistent, and strict wSTL explanations without sacrificing accuracy.
+This work bridges policy learning and explainability through formal methods, contributing to more transparent decision-making in robotics.
 I Introduction
-Recent innovations in learning-based methods, particularly those using neural networks, have significantly improved both high-level decision-making and low-level control in robotics.
-However, these neural network policies are inherently black-box in nature, which poses significant concerns for safety-critical robotic applications, such as autonomous vehicles, where explainability is paramount.
-To address these concerns, various approaches have been explored to enhance explainability of learned policies [1, 2], including post-hoc analysis techniques [3, 4] and methods to learn policies that are inherently explainable [5, 6].
+Recent innovations in learning-based methods, particularly those using neural networks, have improved high-level decision-making and low-level control in robotics.
+However, neural network policies are inherently black-box in nature, which poses significant concerns for safety-critical robotic applications, such as autonomous vehicles, where explainability is paramount.
+Various approaches have been explored to enhance the explainability of learned policies [18], including post-hoc techniques [12] and methods to learn policies that are inherently explainable [14].
 In this paper, we focus on post-hoc analysis to enable explanation of any given robot policy.
-While existing approaches, like saliency maps [7] and feature attribution [8], provide valuable information, they often lack language-based explanations, which limits human-interpretability.
-Autoregressive language models have recently shown promising results for providing language-based explanations [9, 10, 11], but they do not support formal verification of the output explanation, which limits their applicability in domains requiring strong safety and correctness assurances [12].
-Formal methods offer an alternative that emphasizes verifiability.
-In particular, temporal logic has been used to specify robotic tasks and system constraints in an interpretable yet mathematically precise manner that supports verification [13, 14, 15].
-Recent efforts have also used temporal logic, primarily linear temporal logic (LTL) and signal temporal logic (STL), to infer explanations of given trajectory data.
-For example, [16, 17, 18] use a decision tree classifier to infer the LTL or STL specification that best explains the difference between two classes of trajectories.
-[19] uses a greedy explanation search method over explanation candidates in LTL driven by policy similarity metrics.
+While existing approaches, like saliency maps [19], feature attribution [20], and non-temporal concept/rule extraction [23], provide valuable information, they often lack language-based explanations, which limits human interpretability.
+Autoregressive language models have shown promising results for providing language-based explanations [26], but they do not support formal verification of the output explanation, which limits their applicability in safety-critical domains [7].
+Temporal logic offers an alternative approach to specifying robotic tasks and system constraints in an interpretable yet mathematically precise manner that supports verification [8, 4, 5].
+Recent efforts have also used temporal logic to infer explanations of given trajectory data.
+For example, [10, 6] use a decision tree (DT) classifier to infer the linear temporal logic (LTL) or signal temporal logic (STL) specification that best explains the difference between two classes of trajectories.
+[29] uses a greedy search over candidate LTL explanations driven by policy similarity metrics.
 However, these methods are limited in their ability to infer specifications where certain predicates are more critical than others.
-For instance, [18] can infer such relations only for conjunctions.
-Weighted signal temporal logic (wSTL) [20] extends LTL and STL by introducing importance weights over predicates, which can highlight the relative priority of various task components.
+Weighted STL (wSTL) [17] extends STL by introducing importance weights over predicates, which can highlight the relative priority of various task components.
 The expressiveness of wSTL thus makes it particularly suitable for describing robotic behaviors in settings with competing or hierarchical objectives.
-Figure 1: Existing methods for neuro-symbolic post-hoc explainability produce specifications that are verbose (i.e., include unnecessary predicates), inconsistent across training seeds, and loose (e.g., “catch-all” specifications that do not provide useful insights).
-Our proposed network, TLNet, addresses these concerns.
-The presence of real-valued weights in wSTL has enabled the use of neural network approaches for inferring wSTL specifications that explain given trajectory data.
-These methods have shown promising results, where they show high classification performance for various applications, including anomaly detection in naval surveillance [21], fault diagnosis [22], and weather forecasting [23].
+Figure 1: Existing neuro-symbolic methods often produce explanations that are verbose, inconsistent, and loose.
+Our method, TLNet, addresses this by generating concise, consistent, and strict wSTL specifications.
+The presence of real-valued weights in wSTL has enabled the use of neural networks for inferring wSTL specifications that explain given trajectory data.
+These methods have shown high classification accuracy for various applications, including anomaly detection [27], fault diagnosis [25], and weather forecasting [3].
 However, existing work only focuses on classification accuracy, which does not fully capture the actual explainability of learned specifications.
-For example, as shown in Figure 1, these methods often produce specifications that are unnecessarily verbose (i.e., they include many unnecessary predicates), inconsistent across seeds, and/or loose, such that they are satisfied but do not provide useful insights (e.g., “catch-all” specifications that include many conjunctions or overly simplified specifications that do not fully capture the task or constraints).
-Some works address verbosity by incorporating heuristic simplification methods, like only keeping the top-$k$ weighted predicates [21, 24] or using a greedy process to remove weights [25].
-However, as we show in our empirical results, no existing methods effectively balance classification accuracy with verbosity, inconsistency, and looseness.
+For example, as shown in Figure 1, these methods often produce specifications that are unnecessarily verbose (i.e., they include many unnecessary predicates), inconsistent across seeds, and/or loose, such that they are satisfied but do not provide useful insights (e.g., “catch-all” specifications that include many conjunctions or overly simplified specifications).
+Some works address verbosity by incorporating heuristic simplifications, like only keeping the top-$k$ predicates [27, 28] or using a greedy process to remove weights [15].
+However, no existing methods effectively balance classification accuracy with verbosity, inconsistency, and looseness.
 In this work, we propose TLNet, a neural network architecture for inferring wSTL explanations of robot policies that are concise (i.e., include as few predicates as needed), consistent (i.e., produce similar explanations across training seeds), and strict (i.e., contain as few disjunctions as needed).
-Our method targets a subclass of wSTL structures with arbitrary conjunctions and disjunctions of predicates in temporal clauses.
-TLNet includes a multi-step simplification process that reduces specification verbosity and improves consistency, while ensuring strictness is maintained.
-This process includes a predicate filtering step, a regularized loss function, and iterative weight pruning.
-We also provide theoretical justification for our wSTL-based neural network architecture, which essentially reduces learning parameters by the distributivity of importance weight, and introduce novel metrics for quantifying explainability—conciseness, consistency, and strictness—that complement standard classification accuracy, addressing a significant gap in current evaluation practices.
-We demonstrate our method in three simulated robotics environments that vary in continuity, dimensionality, task structure, and policy complexity, and show that it generally outperforms baselines with respect to classification accuracy and our proposed explainability metrics.
-TLNet bridges the gap between neural network-based policy learning and provides formal explanations of learned policies, offering a step toward safer and more transparent robotic systems.
+TLNet achieves this through a multi-step simplification process consisting of predicate filtering, a regularized loss function, and iterative weight pruning.
+We provide theoretical justification for our network’s weight distributivity and introduce explainability metrics—conciseness, consistency, and strictness—to complement standard classification accuracy.
+We demonstrate our method in three simulated robotics environments, showing it outperforms baselines in our proposed metrics while matching or improving classification accuracy.
 II Preliminaries
-Let a trajectory (or a signal) of $n$-dimensional states be a discrete-time series $\tau=s_{0},s_{1},\dots,s_{H}$, where $s_{t}\in S$ is a state at timestep $t\in\mathbb{Z}_{>0}$, $S\subseteq\mathbb{R}^{n}$ is the state space, and $H$ is the time horizon of $\tau$.
+Let a trajectory (or a signal) of $n$-dimensional states be a discrete-time series $\tau=s_{0},s_{1},\dots,s_{H}$, where $s_{t}\in S$ is a state at timestep $t\in\mathbb{Z}_{\geq 0}$, $S\subseteq\mathbb{R}^{n}$ is the state space, and $H$ is the time horizon of $\tau$.
 We also define an interval $I$ as $I:=[a,b]=\{k\in\mathbb{Z}_{\geq 0}|a\leq k\leq b;a,b\in\mathbb{Z}_{\geq 0}\}$, $\lvert I\rvert$ as the cardinality of $I$, and $t+I$ as the interval $[t+a,t+b]$.
@@
 We consider wSTL syntax [17] defined recursively as $\phi:=\top\mid\psi\mid\neg\phi\mid\bigwedge_{i=1}^{N}{}^{w}\phi_{i}\mid\bigvee_{i=1}^{N}{}^{w}\phi_{i}\mid\mathcal{G}_{I}^{w}\phi\mid\mathcal{F}_{I}^{w}\phi$, where $\top$ is the logical True value, $\psi:=f(s)\geq c$ is an atomic predicate given a function $f:S\rightarrow\mathbb{R}$ with constant $c\in\mathbb{R}$, and $\neg$, $\wedge$, and $\vee$ are Boolean negation, conjunction, and disjunction operators, respectively.
+$\mathcal{G}_{I}$ and $\mathcal{F}_{I}$ are globally and eventually operators, respectively.
 For conjunctive and disjunctive operators, we define the weights as $w:=[w_{i}]_{i=1:N}\in\mathbb{R}^{N}_{>0}$, where $N$ is the number of subformulae in the operator.
@@
 We discuss the reasons for this requirement in Section III-E.
-Given this requirement, we define normalized wSTL syntax recursively as $\phi:=\top\mid\psi\mid\neg\phi\mid\bigwedge_{i=1}^{N}{}^{w}\phi_{i}\mid\bigvee_{i=1}^{N}{}^{w}\phi_{i}\mid\mathcal{G}_{I}^{w}\phi\mid\mathcal{F}_{I}^{w}\phi_{i},$ where all terminology carries over from wSTL, except we further require $\sum_{i=1}^{N}w_{i}=\sum_{k\in I}w_{k}=1$, where $w_{i},w_{k}\in[0,1]$.
+Given this requirement, we define normalized wSTL syntax recursively as $\phi:=\top\mid\psi\mid\neg\phi\mid\bigwedge_{i=1}^{N}{}^{w}\phi_{i}\mid\bigvee_{i=1}^{N}{}^{w}\phi_{i}\mid\mathcal{G}_{I}^{w}\phi\mid\mathcal{F}_{I}^{w}\phi,$ where all terminology carries over from wSTL, except we further require $\sum_{i=1}^{N}w_{i}=\sum_{k\in I}w_{k}=1$, where $w_{i},w_{k}\in[0,1]$.
 We allow the weights to be zero because the zero weights do not influence the overall robustness [27].
 To avoid large input values to our neural network, we also define normalized quantitative satisfaction semantics, or robustness, by extending [17] as follows,
-$r(\psi,\tau,t)$ ::=:= $\begin{cases}\frac{f(s_{t})-c}{\sup_{s\in S}f(s)-c}&\text{if $f(s_{t})-c\geq 0$,}\\ \frac{f(s_{t})-c}{\inf_{s\in S}f(s_{t})-c}&\text{otherwise}\end{cases},$ (1)
-$r(\neg\phi,\tau,t)$ ::=:= $-r(\phi,\tau,t),$ (2)
-$r\left(\bigwedge_{i=1}^{N}{}^{w}\phi_{i},\tau,t\right)$ ::=:= $\otimes^{\wedge}(w,[r(\phi_{i},\tau,t)]_{i=1:N}),$ (3)
-$r\left(\bigvee_{i=1}^{N}{}^{w}\phi_{i},\tau,t\right)$ ::=:= $\oplus^{\vee}(w,[r(\phi_{i},\tau,t)]_{i=1:N}),$ (4)
-$r\left(\mathcal{G}^{w}_{I}\phi,\tau,t\right)$ ::=:= $\otimes^{\mathcal{G}}_{I}(w,[r(\phi,\tau,t^{\prime})]_{t^{\prime}\in t+I}),$ (5)
-$r\left(\mathcal{F}^{w}_{I}\phi,\tau,t\right)$ ::=:= $\oplus^{\mathcal{F}}_{I}(w,[r(\phi,\tau,t^{\prime})]_{t^{\prime}\in t+I}),$ (6)
-where the aggregation functions $\otimes^{\wedge}:\mathbb{R}^{N}_{\geq 0}\times\mathbb{R}^{N}\rightarrow\mathbb{R},\otimes^{\vee}:\mathbb{R}^{N}_{\geq 0}\times\mathbb{R}^{N}\rightarrow\mathbb{R},\otimes^{\mathcal{G}}:\mathbb{R}^{\lvert I\rvert}_{\geq 0}\times\mathbb{R}^{\lvert I\rvert}\rightarrow\mathbb{R}$, and $\otimes^{\mathcal{F}}:\mathbb{R}^{\lvert I\rvert}_{\geq 0}\times\mathbb{R}^{\lvert I\rvert}\rightarrow\mathbb{R}$ correspond to the operators $\wedge,\vee,\mathcal{G}$, and $\mathcal{F}$, respectively.
+$r(\psi,\tau,t)$ $:=$ $\begin{cases}\frac{f(s_{t})-c}{\sup_{s\in S}f(s)-c}&\text{if $f(s_{t})-c\geq 0$,}\\ \frac{f(s_{t})-c}{\inf_{s\in S}f(s)-c}&\text{otherwise}\end{cases},$ (1)
+$r(\neg\phi,\tau,t)$ $:=$ $-r(\phi,\tau,t),$ (2)
+$r\left(\bigwedge_{i=1}^{N}{}^{w}\phi_{i},\tau,t\right)$ $:=$ $\otimes^{\wedge}(w,[r(\phi_{i},\tau,t)]_{i=1:N}),$ (3)
+$r\left(\bigvee_{i=1}^{N}{}^{w}\phi_{i},\tau,t\right)$ $:=$ $\oplus^{\vee}(w,[r(\phi_{i},\tau,t)]_{i=1:N}),$ (4)
+$r\left(\mathcal{G}^{w}_{I}\phi,\tau,t\right)$ $:=$ $\otimes^{\mathcal{G}}_{I}(w,[r(\phi,\tau,t^{\prime})]_{t^{\prime}\in t+I}),$ (5)
+$r\left(\mathcal{F}^{w}_{I}\phi,\tau,t\right)$ $:=$ $\oplus^{\mathcal{F}}_{I}(w,[r(\phi,\tau,t^{\prime})]_{t^{\prime}\in t+I}),$ (6)
+where the aggregation functions $\otimes^{\wedge}:\mathbb{R}^{N}_{\geq 0}\times\mathbb{R}^{N}\rightarrow\mathbb{R},\oplus^{\vee}:\mathbb{R}^{N}_{\geq 0}\times\mathbb{R}^{N}\rightarrow\mathbb{R},\otimes^{\mathcal{G}}_{I}:\mathbb{R}^{\lvert I\rvert}_{\geq 0}\times\mathbb{R}^{\lvert I\rvert}\rightarrow\mathbb{R}$, and $\oplus^{\mathcal{F}}_{I}:\mathbb{R}^{\lvert I\rvert}_{\geq 0}\times\mathbb{R}^{\lvert I\rvert}\rightarrow\mathbb{R}$ correspond to the operators $\wedge,\vee,\mathcal{G}$, and $\mathcal{F}$, respectively.
+In practice, the infimum and supremum bounds in Equation 1 can be inferred from physical properties of the environment (we assume known map sizes in our results) or sampled trajectories.
 For notational simplicity, we interchangeably use $r(\phi,\tau,t)$ and $r_{\tau t}^{\phi}$.
@@
 $\oplus^{\vee}(w,[r_{\tau t}^{\phi_{i}}]_{i=1:N})$ $:=\frac{\sum_{i=1}^{N}w_{i}r_{\tau t}^{\phi_{i}}\exp(\frac{r_{\tau t}^{\phi_{i}}}{\sigma})}{\sum_{i=1}^{N}w_{i}\exp(\frac{r_{\tau t}^{\phi_{i}}}{\sigma})},$ (8)
-$\otimes^{\mathcal{G}}(w,[r_{\tau t^{\prime}}^{\phi}]_{t^{\prime}\in t+I})$ $:=\frac{\sum_{t^{\prime}\in t+I}w_{t^{\prime}}r_{\tau t^{\prime}}^{\phi}\exp(\frac{-r_{\tau t^{\prime}}^{\phi}}{\sigma})}{\sum_{t^{\prime}\in t+I}w_{t^{\prime}}\exp(\frac{-r_{\tau t^{\prime}}^{\phi}}{\sigma})},$ (9)
-$\oplus^{\mathcal{F}}(w,[r_{\tau t^{\prime}}^{\phi}]_{t^{\prime}\in t+I})$ $:=\frac{\sum_{t^{\prime}\in t+I}w_{t^{\prime}}r_{\tau t^{\prime}}^{\phi}\exp(\frac{r_{\tau t^{\prime}}^{\phi}}{\sigma})}{\sum_{t^{\prime}\in t+I}w_{t^{\prime}}\exp(\frac{r_{\tau t^{\prime}}^{\phi}}{\sigma})}.$ (10)
-We use $\sigma=0.5$ in this work.
+$\otimes^{\mathcal{G}}_{I}(w,[r_{\tau t^{\prime}}^{\phi}]_{t^{\prime}\in t+I})$ $:=\frac{\sum_{t^{\prime}\in t+I}w_{t^{\prime}}r_{\tau t^{\prime}}^{\phi}\exp(\frac{-r_{\tau t^{\prime}}^{\phi}}{\sigma})}{\sum_{t^{\prime}\in t+I}w_{t^{\prime}}\exp(\frac{-r_{\tau t^{\prime}}^{\phi}}{\sigma})},$ (9)
+$\oplus^{\mathcal{F}}_{I}(w,[r_{\tau t^{\prime}}^{\phi}]_{t^{\prime}\in t+I})$ $:=\frac{\sum_{t^{\prime}\in t+I}w_{t^{\prime}}r_{\tau t^{\prime}}^{\phi}\exp(\frac{r_{\tau t^{\prime}}^{\phi}}{\sigma})}{\sum_{t^{\prime}\in t+I}w_{t^{\prime}}\exp(\frac{r_{\tau t^{\prime}}^{\phi}}{\sigma})}.$ (10)
+We empirically found that decreasing $\sigma$ better approximates the max/min aggregations, but can increase training instability due to larger gradients, while increasing $\sigma$ too much can lead to nonsensical explanations.
 Figure 2: We propose a neuro-symbolic approach for generating an explanation of a given robot policy using wSTL.
@@
 We focus on representing normalized wSTL structures composed of temporal operators containing these predicates in conjunctive normal form (CNF), as follows,
-ϕ=italic-$\phi=$ $0.5\mathcal{F}_{I}^{\bar{w}}\left[\bigwedge_{i=1}^{N_{\mathrm{AP}}}w_{i}^{\mathcal{F}}\bigvee_{j=1}^{2N_{\mathrm{AP}}}w_{ij}^{\mathcal{F}}\psi_{j}\right]$
+$\phi=$ $0.5\mathcal{F}_{I}^{\bar{w}}\left[\bigwedge_{i=1}^{N_{\mathrm{AP}}}w_{i}^{\mathcal{F}}\bigvee_{j=1}^{2N_{\mathrm{AP}}}w_{ij}^{\mathcal{F}}\psi_{j}\right]$
 $\wedge$ $0.5\mathcal{G}_{I}^{\bar{w}}\left[\bigwedge_{i=1}^{N_{\mathrm{AP}}}w_{i}^{\mathcal{G}}\bigvee_{j=1}^{2N_{\mathrm{AP}}}w_{ij}^{\mathcal{G}}\psi_{j}\right],$ (11)
 where $w_{i}^{\mathcal{F}},w_{ij}^{\mathcal{F}},w_{i}^{\mathcal{G}},w_{ij}^{\mathcal{G}}\in[0,1]$ are importance weights and $\psi_{j}\in\Psi\cup\Psi_{\neg}$ are predefined atomic predicates.
-We define the weights $\bar{w}$ for temporal operators over interval $I=[0,H]$ to be $\bar{w}=[\bar{w}_{k}=1/\lvert I\rvert]_{k=1:\lvert I\rvert}$ for trajectory $\tau$.
+We define the weights $\bar{w}$ for temporal operators over interval $I=[0,H]$ to be $\bar{w}=[\bar{w}_{k}=1/\lvert I\rvert]_{k\in I}$.
 We define the conjunction weights of the temporal clauses to be 0.5 to equally weigh the importance of the task ($\mathcal{F}$) and constraint ($\mathcal{G}$) parts of the explanation.
@@
 8:  Prune zeroed and smallest $N_{\mathbf{w}}$ parameters in $\mathbf{w}$.
-9:     Re-normalize $\mathbf{w}$.
-10:     Optimize $\mathbf{w}$ to minimize $\tilde{\mathcal{L}}$.
-11:     if only one non-zero weight in $\mathbf{w}^{\mathcal{F}}$/$\mathbf{w}^{\mathcal{G}}$ then break.
-12:return $\phi$ generated from $\mathbf{w}$ and $\Psi$ using Equation 12.
-We reduce the number of parameters that are optimized by distributing the weights in Equation 11, resulting in a specification of the following form,
-[…]
+9:  Re-normalize and optimize $\mathbf{w}$ to minimize $\tilde{\mathcal{L}}$.
+10:  if only one non-zero weight in $\mathbf{w}^{\mathcal{F}}$$\mathbf{w}^{\mathcal{G}}$ then break.
+11:return $\phi$ generated from $\mathbf{w}$ and $\Psi$ using Equation 12.
+We reduce the number of parameters to be optimized by distributing the weights in Equation 11, resulting in a specification of the following form,
+$\phi_{\mathbf{w}}=$
 $0.5\mathcal{F}_{I}^{\bar{w}}\left[\bigwedge_{i=1}^{N_{\mathrm{AP}}}\bigvee_{j=1}^{2N_{\mathrm{AP}}}\tilde{w}_{ij}^{\mathcal{F}}\psi_{j}\right]\wedge 0.5\mathcal{G}_{I}^{\bar{w}}\left[\bigwedge_{i=1}^{N_{\mathrm{AP}}}\bigvee_{j=1}^{2N_{\mathrm{AP}}}\tilde{w}_{ij}^{\mathcal{G}}\psi_{j}\right],$ (12)
-where $\mathbf{w}=\mathbf{w}^{\mathcal{F}}\mathbin{\|}\mathbf{w}^{\mathcal{G}}$, $\mathbf{w}^{\mathcal{F}}=[\tilde{w}_{ij}^{\mathcal{F}}]_{i=1:N_{\mathrm{AP}},j=1:2N_{\mathrm{AP}}}$, $\mathbf{w}^{\mathcal{G}}=[\tilde{w}_{ij}^{\mathcal{G}}]_{i=1:N_{\mathrm{AP}},j=1:2N_{\mathrm{AP}}}$, $\tilde{w}_{ij}^{\mathcal{F}}=w_{i}^{\mathcal{F}}w_{ij}^{\mathcal{F}}$, and $\tilde{w}_{ij}^{\mathcal{G}}=w_{i}^{\mathcal{F}}w_{ij}^{\mathcal{G}}$.
+where $\mathbf{w}=\mathbf{w}^{\mathcal{F}}\mathbin{\|}\mathbf{w}^{\mathcal{G}}$, $\mathbf{w}^{\mathcal{F}}=[\tilde{w}_{ij}^{\mathcal{F}}]_{i=1:N_{\mathrm{AP}},j=1:2N_{\mathrm{AP}}}$, $\mathbf{w}^{\mathcal{G}}=[\tilde{w}_{ij}^{\mathcal{G}}]_{i=1:N_{\mathrm{AP}},j=1:2N_{\mathrm{AP}}}$, $\tilde{w}_{ij}^{\mathcal{F}}=w_{i}^{\mathcal{F}}w_{ij}^{\mathcal{F}}$, and $\tilde{w}_{ij}^{\mathcal{G}}=w_{i}^{\mathcal{G}}w_{ij}^{\mathcal{G}}$.
+This weight distribution reduces the number of optimized parameters from $2N_{\mathrm{AP}}+4N_{\mathrm{AP}}^{2}$ to $4N_{\mathrm{AP}}^{2}$.
 We justify this weight distribution operation in Section III-E.
-An example TLNet architecture for a dataset containing two predicates is shown in Figure 3.
-Figure 3: Example architecture of the proposed neural network when there are two predicates to consider ($\psi_{1}$ and $\psi_{2}$).
-$\mathbf{w}^{\mathcal{F}}$ and $\mathbf{w}^{\mathcal{G}}$ are weight vectors that are optimized.
-$\phi$ is the inferred explanation to the given trajectory.
+Figure 3 shows a TLNet architecture for a dataset containing two predicates.
+Figure 3: Example TLNet architecture with two predicates, where $\phi_{\mathbf{w}}$ is the explanation generated from $\mathbf{w}$.
 III-B Simplification Process
@@
 $d(\mathcal{D},\psi)=\frac{1}{|\mathcal{D}|}\left[\lvert\mathcal{D}_{1}\rvert,\lvert\mathcal{D}_{2}\rvert,\lvert\mathcal{D}_{3}\rvert\right$], (13)
-$\mathcal{D}_{1}$ :=$:=\set{\tau}{r(\psi,\tau,t)\geq 0\quad\forall t\in[0,H],\tau\in\mathcal{D}},$ (14)
+[…]
 $\mathcal{D}_{2}:=\mathcal{D}\setminus(\mathcal{D}_{1}\cup\mathcal{D}_{3}),$ (15)
-$\mathcal{D}_{3}$ :=$:=\set{\tau}{r(\psi,\tau,t)<0\quad\forall t\in[0,H],\tau\in\mathcal{D},}.$ (16)
+[…]
 That is, $\mathcal{D}_{1}$ is the set of trajectories where $\psi$ is always satisfied, $\mathcal{D}_{2}$ is the set where $\psi$ is satisfied at least once, and $\mathcal{D}_{3}$ is the set where $\psi$ is never satisfied.
 Now, let $\mathcal{T}_{p}=\{(\tau,l)|l=1,(\tau,l)\in\mathcal{T}\}$ be the positive set of trajectories in $\mathcal{T}$ and $\mathcal{T}_{n}=\mathcal{T}\setminus\mathcal{T}_{p}$ be the negative set.
-For each predicate $\psi\in\Psi$, we then calculate $d(\mathcal{T}_{p},\psi)$ and $d(\mathcal{T}_{n},\psi)$, and then calculate the cosine similarity $S$ between those distributions.
-We then remove predicates from the set $\Psi$ (and their negation from $\Psi_{\neg}$) which have a cosine similarity $S>S_{\mathrm{th}}$, where $S_{\mathrm{th}}$ is a user-defined threshold.
+For each predicate $\psi\in\Psi$, we compute $d(\mathcal{T}_{p},\psi)$ and $d(\mathcal{T}_{n},\psi)$, and then calculate the cosine similarity $S$ between those distributions.
+We then remove predicates from the set $\Psi$ (and their negations from $\Psi_{\neg}$) which have a cosine similarity $S\geq S_{\mathrm{th}}$, where $S_{\mathrm{th}}$ is a user-defined threshold.
 III-B2 Loss Regularization
 The regularization step uses two regularizers to improve the conciseness of inferred explanations. The first regularizer, a temporal clause regularizer, $R_{\mathrm{T}}$, is defined as,
-$R_{\mathrm{T}}(\mathbf{w}):=\sum_{j=1}^{N_{\mathrm{AP}}}\max\left(\sum_{i=1}^{N_{\mathrm{AP}}}\tilde{w}_{ij}^{\mathcal{F}},\sum_{i=1}^{N_{\mathrm{AP}}}\tilde{w}_{i(j+N_{\mathrm{AP}})}^{\mathcal{F}}\right)\\ \max\left(\sum_{i=1}^{N_{\mathrm{AP}}}\tilde{w}_{ij}^{\mathcal{G}},\sum_{i=1}^{N_{\mathrm{AP}}}\tilde{w}s_{i(j+N_{\mathrm{AP}})}^{\mathcal{G}}\right).$ (17)
-Intuitively, this regularizer encourages the task and constraint clauses to contain different atomic predicates.
+$R_{\mathrm{T}}(\mathbf{w}):=\sum_{j=1}^{N_{\mathrm{AP}}}\max\left(\sum_{i=1}^{N_{\mathrm{AP}}}\tilde{w}_{ij}^{\mathcal{F}},\sum_{i=1}^{N_{\mathrm{AP}}}\tilde{w}_{i(j+N_{\mathrm{AP}})}^{\mathcal{F}}\right)\\ \max\left(\sum_{i=1}^{N_{\mathrm{AP}}}\tilde{w}_{ij}^{\mathcal{G}},\sum_{i=1}^{N_{\mathrm{AP}}}\tilde{w}_{i(j+N_{\mathrm{AP}})}^{\mathcal{G}}\right).$ (17)
+Intuitively, this regularizer encourages the task and constraint clauses to contain different atomic predicates, and is independent of clause sparsity.
 The second regularizer, a disjunctive clause regularizer, $R_{\mathrm{D}}$, is defined as,
@@
 where $\mathbf{w}^{\mathcal{O}}\in\{\mathbf{w}^{\mathcal{F}},\mathbf{w}^{\mathcal{G}}\}$ defines the weights within one of our temporal operators and $\tilde{w}^{\mathcal{O}}_{ij},\tilde{w}^{\mathcal{O}}_{jk}\in\mathbf{w}^{\mathcal{O}}$.
-Intuitively, this regularizer encourages disjunctive clauses within a temporal clause to contain different predicates.
+Intuitively, this regularizer encourages disjunctive clauses within a temporal clause to contain different predicates, and is again independent of clause sparsity.
 Given $(\tau,l)\in\mathcal{T}$, our overall loss, $\tilde{\mathcal{L}}$, is then,
-~$\tilde{\mathcal{L}}(\tau,l,\mathbf{w})=\mathcal{L}(\tau,l,\mathbf{w})+$ $\lambda_{R_{\mathrm{T}}}R_{\mathrm{T}}(\mathbf{w})$
+$\tilde{\mathcal{L}}(\tau,l,\mathbf{w})=\mathcal{L}(\tau,l,\mathbf{w})+$ $\lambda_{R_{\mathrm{T}}}R_{\mathrm{T}}(\mathbf{w})$
 $+$ $\lambda_{R_{\mathrm{D}}}[R_{\mathrm{D}}(\mathbf{w}^{\mathcal{F}})+R_{\mathrm{D}}(\mathbf{w}^{\mathcal{G}})],$ (19)
-where $\mathcal{L}$ is a classification loss and $\lambda_{R_{\mathrm{T}}},\lambda_{R_{\mathrm{D}}}\in\mathbb{R}_{\geq 0}$ are regularizer hyperparameters.
-We use $\lambda_{R_{\mathrm{T}}}=0.01$ and $\lambda_{R_{\mathrm{D}}}=0.1$ for this work.
-We use the classification loss from [21],
+where $\mathcal{L}$ is a classification loss and $\lambda_{R_{\mathrm{T}}},\lambda_{R_{\mathrm{D}}}\in\mathbb{R}_{\geq 0}$ are regularizer hyperparameters. We use the classification loss from [27],
 $\mathcal{L}(\tau,l,\mathbf{w})=\exp(-\zeta lr(\phi_{\mathbf{w}},\tau)),$ (20)
-where $\zeta\in\mathbb{R}_{>0}$ is a hyperparameter ($\zeta=1$ for this work).
+where $\zeta\in\mathbb{R}_{>0}$ is a hyperparameter.
 III-B3 Weight Pruning
@@
 The remaining weights are then normalized.
-The pruning process terminates if there is only one non-zero weight in $\mathbf{w}^{\mathcal{F}}$ or $\mathbf{w}^{\mathcal{G}}$ to ensure the task-constraint structure of the inferred explanation.
+The pruning process is performed $N_{\mathrm{pr}}\in\mathbb{Z}_{>0}$ times and terminates if there is only one non-zero weight in $\mathbf{w}^{\mathcal{F}}$ or $\mathbf{w}^{\mathcal{G}}$ to ensure the desired task-constraint structure.
+Empirically, increasing $N_{\mathrm{w}}$ leads to training instability, since pruning changes the computation graph structure of the neural network.
+Decreasing $N_{\mathrm{pr}}$ can lead to unnecessarily complex explanations.
 III-C Trajectory Sampling
@@
 III-D Explainability Evaluation Metrics
-Beyond classification performance, TLNet also emphasizes the human-interpretability of inferred explanations.
+Beyond classification performance, TLNet also emphasizes the human interpretability of inferred explanations.
 We quantify this emphasis by introducing three explainability metrics—conciseness, consistency, and strictness.
@@
 $\mathrm{Consistency}:=$
-$\frac{1}{N}\sum_{n=1}^{N}\begin{cases}\max_{\sigma_{n}\in\Sigma_{n}}\sum_{\hat{\phi}_{n}\in\hat{\Phi}_{n}}\frac{\mathbf{1}_{\sigma_{n}}(\hat{\phi}_{n})}{K\lvert\Sigma_{n}\rvert}&\text{if $\Sigma_{n}\neq\emptyset$},\\ 0&\text{otherwise},\end{cases}$
-where $\mathbf{1}_{\sigma_{n}}(\cdot)$ is the indicator function for $\sigma_{n}$.
+$\frac{1}{N}\sum_{n=1}^{N}\begin{cases}\max_{\sigma_{n}\in\Sigma_{n}}\sum_{\hat{\phi}_{n}\in\hat{\Phi}_{n}}\frac{\mathbb{I}(\sigma_{n}=\hat{\phi}_{n})}{K\lvert\Sigma_{n}\rvert}&\text{if $\Sigma_{n}\neq\emptyset$},\\ 0&\text{otherwise}.\end{cases}$
 Consistency ranges from 0 to 1, with higher scores indicating greater agreement in clause structures across explanations, reflecting structural stability in learned explanations.
@@
 Strictness ranges from 0 to 1, with higher scores implying stricter logical expressions by rewarding conjunctions and penalizing disjunctions.
+For example, given a temporal clause $\phi_{0}=\mathcal{F}[(0.5\psi_{0}\vee 0.25\psi_{1})\wedge 0.25\psi_{2}]$ and $P=6$, we get $C_{0}=1$ and $D_{0}=1$, with strictness of 0.167 for $\phi_{0}$.
 III-E Weight Distribution in wSTL
@@
 The normalized weights of a wSTL formula in CNF can be syntactically distributed across Boolean operators as follows.
+[^1] Syntactic weight distributivity also holds for temporal operators by distributing operator weights to the normalized interval weights followed by distributions to the inner Boolean structures.
 Theorem 1 (Syntactic Weight Distributivity of Boolean Operators).
@@
 Proof.
-[…]
+$\{\text{LHS}\}$ $=w_{1}(w_{11}(\neg)\psi_{11}\vee...\vee w_{1m_{1}}(\neg)\psi_{1m_{1}})$
 ... $\wedge w_{n}(w_{n1}(\neg)\psi_{n1}\vee...\vee w_{nm_{n}}(\neg)\psi_{nm_{n}})$
@@
 We perform a series of empirical experiments to compare our method to relevant baselines for robotic tasks.
-We consider three baselines for inference of wSTL explanations, each adapted to model our normalized wSTL structures (given by Equation 12): {outline} \1 NN-TLI [25]: this method implements a greedy pruning strategy that independently removes network weights that do not affect classification accuracy after optimization.
-We removed their weight-rounding process (to convert to STL) to consider the structure in Equation 12 and used the aggregation functions given in Figure 2, since our empirical results showed more training stability with those functions.
-\1 STONE (top-3) [21]: this method retains the 3 largest network weights after training.
-\1 STONE (top-5) [21, 24]: this method retains the 5 largest network weights after training.
-All baselines were implemented using the same neural network architecture as TLNet to ensure a consistent basis for comparison.
+We consider three baselines for inference of wSTL explanations, each adapted to model our normalized wSTL structures (given by Equation 12):
+  • NN-TLI [15]: this method implements a greedy pruning strategy that independently removes network weights that do not affect classification accuracy after optimization.
+We removed their weight-rounding process (to convert to STL) to consider the structure in Equation 12 and used the aggregation functions given in Equations 7, 8, 9 and 10, since our empirical results showed more training stability with those functions.
+  • STONE (top-3, top-5) [27, 28]: this method retains the 3 or 5 largest network weights after training.
+  • DT (STL) [6]: this method uses decision trees to mine STL (not wSTL) explanations.
+The first two baselines, NN-TLI and STONE (top-3, top-5), were implemented using the same neural network architecture as TLNet to ensure a consistent basis for comparison.
 For each policy we explain, we collected 500 positive trajectories using the trained policy and 500 negative trajectories using a random-walk policy.
 Each dataset was split 80% and 20% for training and evaluation, respectively.
-Our code is available on GitHub after acceptance.
+Hyperparameter values are summarized in Table I.
+Our code, including videos of target policies, is available online [^2] https://github.com/LIRA-illinois/tlnet.
 (a) CtF environment.
@@
 (c) Robot navigation.
-Figure 4: Screenshots from our environments.
+Figure 4: Initial states of our test environments with relevant objects shown.
+TABLE I: Hyperparameter values and their ranges.
+Environments $\sigma$ $\zeta$ $S_{\mathrm{th}}$ $\lambda_{R_{\mathrm{T}}}$ $\lambda_{R_{\mathrm{D}}}$ $N_{\mathrm{pr}}$ $N_{\mathrm{w}}$ $H$
+CtF 0.5 1.0 0.99 0.01 0.1 34 1 100
+Fetch Push 0.5 1.0 0.99 0.01 0.1 20 1 200
+Robot Nav. 0.5 1.0 0.99 0.01 0.1 28 1 400
+Range ($\in$) $\mathbb{R}_{>0}$ $\mathbb{R}_{>0}$ [-1, 1] $\mathbb{R}_{\geq 0}$ $\mathbb{R}_{\geq 0}$ $\mathbb{Z}_{>0}$ $\mathbb{Z}_{>0}$ $\mathbb{Z}_{>0}$
 IV-A Test Environments
@@
 We considered blue agent policies optimized for five scenarios.
-The first four considered different red agent heuristic policies (capture, fight, patrol, mixture) with the same reward function (capture: +1, kill: +0.5, being captured: $-1$, timestep penalty: $-0.01$).
+The first four considered different red agent heuristic policies (capture, fight, patrol, mix) with the same reward function (capture: +1, kill: +0.5, being captured: 1-1, timestep penalty: 0.01-0.01).
 The fifth scenario (capture 0) used a capture red agent with the reward function modified to remove the kill reward (kill: 0).
-The red agent policies are defined as follows: capture pursues the flag; fight pursues the blue agent; patrol defends a boundary region; and mixture switches between capturing and defending.
+The red agent policies are defined as follows: capture pursues the flag; fight pursues the blue agent; patrol defends a boundary region; and mix uses fight when the blue agent is in the red territory and capture otherwise.
 Red agents perform a random action 25% of the time.
@@
 We also consider a continuous control task based on the Fetch mobile manipulator [9].
-The robot must push a blue block (“b”) to a red target (“t”) using its gripper (“g”), without disturbing a black obstacle (“o”).
+The robot must push a blue block (“b”) to a red target (“t”) using its grip (“g”), without disturbing a black obstacle (“o”).
 The state and action spaces are continuous ($s\in\mathbb{R}^{55},a\in\mathbb{R}^{4}$).
-Eight predicates were defined based on the distances between the gripper, block, target, and obstacle, including drop-related (“d”) conditions: $\psi_{\mathrm{gb}}$, $\psi_{\mathrm{gt}}$, $\psi_{\mathrm{go}}$, $\psi_{\mathrm{bt}}$, $\psi_{\mathrm{bo}}$, $\psi_{\mathrm{ot}}$, $\psi_{\mathrm{bd}}$, $\psi_{\mathrm{od}}$.
+Eight predicates were defined based on the distances between the grip, block, target, and obstacle, including drop-related (“d”) conditions: $\psi_{\mathrm{gb}}$, $\psi_{\mathrm{gt}}$, $\psi_{\mathrm{go}}$, $\psi_{\mathrm{bt}}$, $\psi_{\mathrm{bo}}$, $\psi_{\mathrm{ot}}$, $\psi_{\mathrm{bd}}$, $\psi_{\mathrm{od}}$.
 IV-A3 Robot Navigation
@@
 We included predicates irrelevant to task success in each environment (e.g., $\psi_{\mathrm{ra,ob}}$, $\psi_{\mathrm{ot}}$, $\psi_{\mathrm{hv}}$) to evaluate whether the considered methods could eliminate them from their explanations.
+All predicates are distance-based (e.g., $d_{\mathrm{ba,ra}}<0.5$ ) except for $\psi_{\mathrm{ra,df}}$ (binary).
+Average TLNet optimization times were 170.8 s (CtF), 118.3 s (Fetch Push), and 116.1 s (Robot Navigation).
+[^3] With an Intel Xeon Silver 4314 CPU and NVIDIA RTX A4000 GPU.
 IV-B Policy Optimization
@@
 We used Stable-Baselines3 [21] for training all policies.
-For the CtF and navigation environments, PPO [30] was used; for the Fetch Push task, we used SAC [31] with hindsight experience replay (HER) [32].
+For the CtF and navigation environments, we used PPO [24]; for the Fetch Push task, we used SAC [11] with hindsight experience replay (HER) [2].
 Hyperparameters were taken from prior work [29, 22, 13].
 For each policy, we performed 10 independent optimization runs with random initialization and collected classification accuracy and explainability metrics.
-The time horizon ($H$) for each environment is 100 for CtF, 200 for Fetch Push, and 400 for Robot Navigation.
-TABLE I: Comparison of the Most Frequently Inferred Explanations.
-Scenarios TLNet (Ours) NN-TLI STONE (top-3) STONE (top-5)
-CtF Capture
+TABLE II: Comparison of the Most Frequently Inferred Explanations with Frequencies (%).
+Scenario TLNet (Ours) NN-TLI STONE (top-3) STONE (top-5) DT
+CtF
+Capture
 $0.5\mathcal{F}_{I}^{\bar{w}}[1.0\psi_{\mathrm{ba,rf}}]\wedge$
-$0.5\mathcal{G}_{I}^{\bar{w}}[0.30\psi_{\mathrm{ba,rf}}\vee 0.70\neg\psi_{\mathrm{ra,bf}}]$
-$0.5\mathcal{F}_{I}^{\bar{w}}[0.26\psi_{\mathrm{ba,rf}}\wedge(0.25\psi_{\mathrm{ba,rf}}\vee 0.33\neg\psi_{\mathrm{ra,bt}})\wedge$
-$(0.09\neg\psi_{\mathrm{ba,bt}}\vee 0.07\neg\psi_{\mathrm{ra,bt}})]\wedge$
+$0.5\mathcal{G}_{I}^{\bar{w}}[0.30\psi_{\mathrm{ba,rf}}\vee$
+   $0.70\neg\psi_{\mathrm{ra,bf}}]$
+60%
+$0.5\mathcal{F}_{I}^{\bar{w}}[0.26\psi_{\mathrm{ba,rf}}\wedge(0.25\psi_{\mathrm{ba,rf}}\vee 0.33$
+$\neg\psi_{\mathrm{ra,bt}})\wedge(0.09\neg\psi_{\mathrm{ba,bt}}\vee 0.07\neg\psi_{\mathrm{ra,bt}})]\wedge$
 $0.5\mathcal{G}_{I}^{\bar{w}}[0.36\psi_{\mathrm{ba,rf}}\vee 0.64\neg\psi_{\mathrm{ra,bf}}]$
-$\mathcal{F}_{I}^{\bar{w}}[0.73\neg\psi_{\mathrm{ra,bt}}$
-$\wedge 0.27\neg\psi_{\mathrm{ba,bt}}]$
-$\mathcal{F}_{I}^{\bar{w}}[0.83\neg\psi_{\mathrm{ra,bt}}$
-$\wedge 0.17\psi_{\mathrm{ba,rf}}]$
-CtF Fight
+20%
+$\mathcal{F}_{I}^{\bar{w}}[0.73\neg$
+$\psi_{\mathrm{ra,bt}}\wedge 0.$
+$27\neg\psi_{\mathrm{ba,bt}}]$
+50%
+$\mathcal{F}_{I}^{\bar{w}}[0.83\neg$
+$\psi_{\mathrm{ra,bt}}\wedge 0$
+$17\psi_{\mathrm{ba,rf}}]$
+60%
+$\mathcal{F}_{I}[\psi_{\mathrm{ba,rf}}$
+$\vee\psi_{\mathrm{ra,bf}}\vee$
+$\neg\psi_{\mathrm{ba,bt}}]$
+100%
+CtF
+Fight
 $0.5\mathcal{F}_{I}^{\bar{w}}[1.0\psi_{\mathrm{ba,rf}}]\wedge$
-$0.5\mathcal{G}_{I}^{\bar{w}}[0.33\psi_{\mathrm{ba,rf}}\vee 0.67\neg\psi_{\mathrm{ba,ra}}]$
+$0.5\mathcal{G}_{I}^{\bar{w}}[0.33\psi_{\mathrm{ba,rf}}\vee$
+   $0.67\neg\psi_{\mathrm{ba,ra}}]$
+70%
 $0.5\mathcal{F}_{I}^{\bar{w}}[0.77\psi_{\mathrm{ba,rf}}\wedge(0.15\psi_{\mathrm{ba,rf}}\vee 0.08\psi_{\mathrm{ra,bf}})]\wedge$
@@
 $0.16\neg\psi_{\mathrm{ba,bt}}\vee 0.11\neg\psi_{\mathrm{ra,df}})]$
-$\mathcal{F}_{I}^{\bar{w}}[1.0\psi_{\mathrm{ba,rf}}]$ $\mathcal{F}_{I}^{\bar{w}}[1.0\psi_{\mathrm{ba,rf}}]$
-CtF Patrol
+70%
+$\mathcal{F}_{I}^{\bar{w}}$
+$1.0\psi_{\mathrm{ba,rf}}]$
+100%
+$\mathcal{F}_{I}^{\bar{w}}$
+$1.0\psi_{\mathrm{ba,rf}}]$
+100%
+$\mathcal{F}_{I}[\psi_{\mathrm{ba,rf}}]$
+$\wedge$
+$\mathcal{G}_{I}[\psi_{\mathrm{ba,bt}}]$
+100%
+CtF
+Patrol
 $0.5\mathcal{F}_{I}^{\bar{w}}[1.0\psi_{\mathrm{ba,rf}}]\wedge$
-$0.5\mathcal{G}_{I}^{\bar{w}}[0.52\psi_{\mathrm{ba,rf}}\vee 0.48\neg\psi_{\mathrm{ra,bt}}]$
+$0.5\mathcal{G}_{I}^{\bar{w}}[0.52\psi_{\mathrm{ba,rf}}\vee$
+   $0.48\neg\psi_{\mathrm{ra,bt}}]$
+70%
 $0.5\mathcal{F}_{I}^{\bar{w}}[0.35\psi_{\mathrm{ba,rf}}\wedge(0.09\psi_{\mathrm{ba,rf}}\vee 0.55\psi_{\mathrm{ra,df}})]\wedge$
@@
 $0.04\psi_{\mathrm{ra,bf}}\wedge 0.05\neg\psi_{\mathrm{ra,bf}}]$
-$\mathcal{F}_{I}^{\bar{w}}[1.0\psi_{\mathrm{ra,df}}]$ $\mathcal{F}_{I}^{\bar{w}}[1.0\psi_{\mathrm{ra,df}}]$
-CtF Mixture
+30%
+$\mathcal{F}_{I}^{\bar{w}}$
+$1.0\psi_{\mathrm{ra,df}}]$
+100%
+$\mathcal{F}_{I}^{\bar{w}}$
+$1.0\psi_{\mathrm{ra,df}}]$
+90%
+$\mathcal{G}_{I}[\neg\psi_{\mathrm{ba,rf}}$
+$\wedge\psi_{\mathrm{ba,bt}}]$
+100%
+CtF
+Mix
 $0.5\mathcal{F}_{I}^{\bar{w}}[1.0\psi_{\mathrm{ba,rf}}]\wedge$
-$0.5\mathcal{G}_{I}^{\bar{w}}[0.44\psi_{\mathrm{ba,rf}}\vee 0.56\neg\psi_{\mathrm{ra,bf}}]$
-$0.5\mathcal{F}_{I}^{\bar{w}}[1.0\psi_{\mathrm{ba,rf}}]\wedge 0.5\mathcal{G}_{I}^{\bar{w}}[0.21\neg\psi_{\mathrm{ra,bf}}\wedge 0.47\psi_{\mathrm{ba,rf}}$
-$\wedge(0.04\neg\psi_{\mathrm{ra,bf}}\vee 0.05\neg\psi_{\mathrm{ra,df}})\wedge(0.18\psi_{\mathrm{ba,rf}}\vee$
-$0.03\psi_{\mathrm{ra,df}}\vee 0.04\neg\psi_{\mathrm{ra,bf}}\vee 0.02\neg\psi_{\mathrm{ra,bt}})]$
-$\mathcal{F}_{I}^{\bar{w}}[1.0\psi_{\mathrm{ba,bt}}]$
-$\mathcal{F}_{I}^{\bar{w}}[0.48\neg\psi_{\mathrm{ba,bt}}$
-$\wedge 0.52\psi_{\mathrm{ba,rf}}]$
-CtF Capture 0
+$0.5\mathcal{G}_{I}^{\bar{w}}[0.44\psi_{\mathrm{ba,rf}}\vee$
+   $0.56\neg\psi_{\mathrm{ra,bf}}]$
+60%
+$0.5\mathcal{F}_{I}^{\bar{w}}[1.0\psi_{\mathrm{ba,rf}}]\wedge 0.5\mathcal{G}_{I}^{\bar{w}}[0.21\neg\psi_{\mathrm{ra,bf}}\wedge 0.47$
+$\psi_{\mathrm{ba,rf}}\wedge(0.04\neg\psi_{\mathrm{ra,bf}}\vee 0.05\neg\psi_{\mathrm{ra,df}})\wedge(0.18\psi$
+$_{\mathrm{ba,rf}}\vee 0.03\psi_{\mathrm{ra,df}}\vee 0.04\neg\psi_{\mathrm{ra,bf}}\vee 0.02\neg\psi_{\mathrm{ra,bt}})]$
+10%
+$\mathcal{F}_{I}^{\bar{w}}$
+$1.0\psi_{\mathrm{ba,bt}}]$
+60%
+$\mathcal{F}_{I}^{\bar{w}}[0.48\neg$
+$\psi_{\mathrm{ba,bt}}\wedge 0$
+$52\psi_{\mathrm{ba,rf}}]$
+40%
+$\mathcal{F}_{I}[\neg\psi_{\mathrm{ba,rf}}$
+$\vee\neg\psi_{\mathrm{ba,ra}}]$
+$\wedge\mathcal{G}_{I}[\psi_{\mathrm{ba,rf}}]$
+100%
+CtF
+Capture
+0
 $0.5\mathcal{F}_{I}^{\bar{w}}[1.0\psi_{\mathrm{ba,rf}}]\wedge$
-$0.5\mathcal{G}_{I}^{\bar{w}}[0.31\psi_{\mathrm{ba,rf}}\vee 0.69\neg\psi_{\mathrm{ra,bf}}]$
+$0.5\mathcal{G}_{I}^{\bar{w}}[0.31\psi_{\mathrm{ba,rf}}\vee$
+   $0.69\neg\psi_{\mathrm{ra,bf}}]$
+80%
 $0.5\mathcal{F}_{I}^{\bar{w}}[0.08\psi_{\mathrm{ba,rf}}\wedge(0.40\psi_{\mathrm{ba,rf}}\vee 0.31\neg\psi_{\mathrm{ra,bt}})\wedge$
-$(0.07\neg\psi_{\mathrm{ba,bt}}\vee 0.14\neg\psi_{\mathrm{ra,bt}})]\wedge 0.5\mathcal{G}_{I}^{\bar{w}}[(0.33\psi_{\mathrm{ba,rf}}$
-$\vee 0.58\neg\psi_{\mathrm{ra,bf}})\wedge(0.05\psi_{\mathrm{ba,rf}}\vee 0.04\neg\psi_{\mathrm{ra,bt}})]$
-$\mathcal{F}_{I}^{\bar{w}}[0.67\neg\psi_{\mathrm{ra,bt}}$
-$\wedge 0.33\neg\psi_{\mathrm{ba,bt}}]$
-$\mathcal{F}_{I}^{\bar{w}}[0.83\neg\psi_{\mathrm{ra,bt}}$
-$\wedge 0.17\psi_{\mathrm{ba,rf}}]$
-Fetch Push
-$0.5\mathcal{F}_{I}^{\bar{w}}[1.0\psi_{\mathrm{bt}}]\wedge$
-$0.5\mathcal{G}_{I}^{\bar{w}}[0.41\psi_{\mathrm{gb}}\vee 0.59\psi_{\mathrm{bt}}]$
+$(0.07\neg\psi_{\mathrm{ba,bt}}\vee 0.14\neg\psi_{\mathrm{ra,bt}})]\wedge 0.5\mathcal{G}_{I}^{\bar{w}}[(0.33\psi_{\mathrm{ba,}}$
+$_{\mathrm{rf}}\vee 0.58\neg\psi_{\mathrm{ra,bf}})\wedge(0.05\psi_{\mathrm{ba,rf}}\vee 0.04\neg\psi_{\mathrm{ra,bt}})]$
+20%
+$\mathcal{F}_{I}^{\bar{w}}[0.67\neg$
+$\psi_{\mathrm{ra,bt}}\wedge 0.$
+$33\neg\psi_{\mathrm{ba,bt}}]$
+60%
+$\mathcal{F}_{I}^{\bar{w}}[0.83\neg$
+$\psi_{\mathrm{ra,bt}}\wedge 0$
+$17\psi_{\mathrm{ba,rf}}]$
+40%
+$\mathcal{F}_{I}[\psi_{\mathrm{ba,rf}}\vee$
+$\neg\psi_{\mathrm{ba,ra}}\vee$
+$\neg\psi_{\mathrm{ba,bt}}]$
+100%
+Fetch
+Push
+$0.5\mathcal{F}_{I}^{\bar{w}}[1.0\psi_{\mathrm{bt}}]\wedge 0.5$
+$\mathcal{G}_{I}^{\bar{w}}[0.41\psi_{\mathrm{gb}}\vee 0.59\psi_{\mathrm{bt}}]$
+100%
 $0.5\mathcal{F}_{I}^{\bar{w}}[0.74\psi_{\mathrm{bt}}\wedge(0.10\psi_{\mathrm{bt}}\vee 0.16\psi_{\mathrm{od}})]\wedge$
 $0.5\mathcal{G}_{I}^{\bar{w}}[1.0\neg\psi_{\mathrm{bd}}]$
-$\mathcal{G}_{I}^{\bar{w}}[1.0\psi_{\mathrm{bt}}]$ $\mathcal{G}_{I}^{\bar{w}}[1.0\psi_{\mathrm{bt}}]$
-Robot Navi.
-$0.5\mathcal{F}_{I}^{\bar{w}}[1.0\psi_{\mathrm{eg}}]\wedge 0.5\mathcal{G}_{I}^{\bar{w}}[1.0\neg\psi_{\mathrm{ec}}]$ $\mathcal{F}_{I}^{\bar{w}}[1.0\psi_{\mathrm{eg}}]$ $\mathcal{F}_{I}^{\bar{w}}[1.0\psi_{\mathrm{eg}}]$ $\mathcal{F}_{I}^{\bar{w}}[1.0\psi_{\mathrm{eg}}]$
-TABLE II: Comparison of Conciseness, Consistency, Strictness, and Accuracy Metrics. The best results are bolded.
-Conciseness Consistency Strictness Accuracy
-STONE STONE STONE STONE
-Scenarios Ours NN-TLI Top-3 Top-5 Ours NN-TLI Top-3 Top-5 Ours NN-TLI Top-3 Top-5 Ours NN-TLI Top-3 Top-5
-CtF Capture 0.41 0.21 0.20 0.22 0.33 0.08 0.08 0.13 0.33 0.08 0.08 0.13 0.95 0.95 0.95 0.94
-CtF Fight 0.39 0.18 0.25 0.25 0.68 0.09 0.50 0.50 0.68 0.09 0.50 0.50 0.94 0.92 0.94 0.93
-CtF Patrol 0.38 0.18 0.25 0.28 0.63 0.05 0.50 0.53 0.63 0.05 0.50 0.53 0.95 0.93 0.94 0.94
-CtF Mixture 0.39 0.12 0.21 0.16 0.27 0.06 0.10 0.09 0.27 0.06 0.10 0.09 0.87 0.89 0.88 0.88
-CtF Capt. 0 0.42 0.21 0.20 0.30 0.45 0.11 0.15 0.24 0.45 0.11 0.15 0.24 0.95 0.95 0.95 0.95
-Fetch Push 0.42 0.31 0.25 0.25 1.00 0.28 0.50 0.50 1.00 0.28 0.50 0.50 0.99 0.97 0.97 0.97
-Robot Navi. 0.48 0.22 0.25 0.25 0.68 0.15 0.50 0.50 0.68 0.15 0.50 0.50 0.97 0.96 0.96 0.96
+40%
+$\mathcal{G}_{I}^{\bar{w}}[1.0\psi_{\mathrm{bt}}]$
+100%
+$\mathcal{G}_{I}^{\bar{w}}[1.0\psi_{\mathrm{bt}}]$
+100% $\mathcal{G}_{I}[\neg\psi_{\mathrm{og}}]$ 100%
+Robot
+Nav.
+$0.5\mathcal{F}_{I}^{\bar{w}}[1.0\psi_{\mathrm{eg}}]\wedge$
+$0.5\mathcal{G}_{I}^{\bar{w}}[1.0\neg\psi_{\mathrm{ec}}]$
+70% $\mathcal{F}_{I}^{\bar{w}}[1.0\psi_{\mathrm{eg}}]$ 70%
+$\mathcal{F}_{I}^{\bar{w}}[1.0\psi_{\mathrm{eg}}]$
+100%
+$\mathcal{F}_{I}^{\bar{w}}[1.0\psi_{\mathrm{eg}}]$
+100%
+$\mathcal{F}_{I}[\psi_{\mathrm{eg}}\vee$
+$\psi_{\mathrm{ec}}\vee\psi_{\mathrm{eh}}]$
+100%
+TABLE III: Comparison of Conciseness, Consistency, Strictness with standard deviation.
+The best results are bolded.
+[^5] Statistical significance (Friedman test, Dunn-Bonferroni analysis) against the second best values indicated with * ($p<0.05$), ** ($p<0.01$).
+Conciseness Consistency Strictness
+STONE STONE STONE
+Scenario Ours NN-TLI top-3 top-5 DT Ours NN-TLI top-3 top-5 DT Ours NN-TLI top-3 top-5 DT
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+TABLE IV: Comparison of Accuracy. The best results are bolded.
+Accuracy
+STONE
+Scenario Ours NN-TLI top-3 top-5 DT
+CtF Capt. $\textbf{0.95}_{\pm 0.02}$ $\textbf{0.95}_{\pm 0.01}$ $\textbf{0.95}_{\pm 0.01}$ $0.94_{\pm 0.02}$ 0.95±0.00
+CtF Fight 0.94±0.02 0.92±0.02 0.94±0.02 0.93±0.01 0.94±0.00
+CtF Patrol 0.95±0.02 0.93±0.02 0.94±0.01 0.94±0.02 0.95±0.00
+CtF Mix 0.87±0.01 0.89±0.04 0.88±0.03 0.88±0.05 0.89±0.00
+CtF Capt.0 0.95±0.02 0.95±0.01 0.95±0.01 0.95±0.02 0.95±0.00
+[…]
+Robot Nav. 0.97±0.01 0.96±0.01 0.96±0.01 0.96±0.01 0.98±0.01
 IV-C Experimental Results
-IV-C1 Benefits and Limitations of TLNet
-Table I presents the most frequent explanation obtained by each method for each scenario.
-Across all scenarios, TLNet consistently inferred both task and constraint clauses, while also producing concise and relatively easy-to-understand explanations.
+IV-C1 Benefits & Limitations of TLNet
+Table II presents the most frequent explanation obtained by each method for each scenario.
+Across all scenarios, TLNet consistently inferred both task and constraint clauses, while also producing concise and relatively comprehensible explanations.
 In comparison, NN-TLI typically yielded verbose explanations that are difficult to understand, while STONE often captured only one temporal clause, resulting in incomplete explanations.
-The explanations produced by TLNet are also qualitatively reasonable (and often correct).
-For example, for all CtF scenarios, TLNet correctly inferred the task clause as $\psi_{\mathrm{ba,rf}}$ (“blue agent captures the red flag”)—no other method correctly inferred this task for every scenario.
+The explanations produced by TLNet are also qualitatively reasonable.
+For all CtF scenarios, TLNet correctly inferred the task clause as $\psi_{\mathrm{ba,rf}}$ (“blue agent captures the red flag”)—no other method correctly inferred this task for every scenario.
 The inferred CtF constraints are also syntactically reasonable.
-For example, in the Capture scenario, the blue agent (global) constraint includes “red agent not on the blue flag” because the capture red policy only targets the blue flag, which required the blue agent to learn to prevent the red agent from reaching the blue flag.
+For example, in the Capture scenario, the blue agent (global) constraint includes “red agent not on the blue flag” because the capture red policy only targets the blue flag, which requires the blue agent to learn to prevent the red agent from reaching the blue flag.
 Other scenarios, such as Fight and Patrol, did not include this predicate in the inferred constraint because those red policies do not target the blue flag.
 In comparison, the CtF constraints inferred by NN-TLI are generally too verbose to understand, while STONE did not infer any constraints.
+The DT (STL) baseline often produced a task- or constraint-only STL explanation, typically with semantically-inaccurate clauses.
 For Fetch Push, TLNet also correctly inferred the task clause as $\psi_{\mathrm{bt}}$ (“block reaches the target”), which no other method correctly inferred.
 While TLNet inferred part of the expected constraint ($\psi_{\mathrm{gb}},\psi_{\mathrm{bt}}$), it missed an expected constraint $\neg\psi_{\mathrm{go}}$ (“grip does not touch the obstacle”), likely due to insufficient occurrences of the grip touching the obstacle in the sampled data.
-Future research could mitigate this issue by introducing more intelligent sampling strategies for generating the negative dataset (e.g., adversarial or targeted trajectory sampling).
 NN-TLI also inferred part of the expected constraint, but also missed the $\neg\psi_{\mathrm{go}}$ constraint.
@@
 For Robot Navigation, all methods correctly inferred the task clause but TLNet was the only one to infer the correct constraint clause as $\neg\psi_{\mathrm{ec}}$ (“the robot does not hit the chaser”).
-As shown in Table II, TLNet also outperformed all baselines in conciseness, consistency, and strictness, with marginal differences in classification accuracy.
-Across all scenarios, TLNet achieved up to 1.9$\times$ higher conciseness, 2.6$\times$ higher consistency, and 2.7$\times$ higher strictness.
-These results support the qualitative findings above, validating the explainability of our method’s outputs and the use of these explainability metrics.
+[…]
 IV-C2 Benefit of Including Weights
-We now demonstrate the benefit of including weights within inferred explanations (e.g., using wSTL rather than STL [25] or LTL [16, 17, 18]) by examining how inferred weights differ across CtF scenarios.
-Consider the Capture and Mixture explanations in Table I.
-Both capture and mixture red policies target the blue flag, but the capture policy is more aggressive in doing so.
-The blue agent, in response, learns slightly different policies: a more defensive policy in Capture, prioritizing the protection of its flag, and a more aggressive policy in Mixture, focusing more on capturing the red flag.
-This policy difference is reflected in the weights in the constraint clause, where the $\neg\psi_{\mathrm{ra,bf}}$ predicate receives a higher weight in the Capture explanation (0.70) than in the Mixture explanation (0.56), due to the Capture policy’s prioritization of flag protection.
-Conversely, the Mixture policy shows a higher weight for the $\psi_{\mathrm{ba,rf}}$ predicate (0.44) than the Capture policy (0.30), due to the Mixture policy’s prioritization of flag capture.
+We now demonstrate the benefit of including weights within inferred explanations (e.g., using wSTL rather than STL [15, 6, 1] or LTL [10]) by examining how inferred weights differ across CtF scenarios.
+Consider the Capture and Mix explanations in Table II.
+Both capture and mix red policies target the blue flag, but the capture policy is more aggressive in doing so.
+The blue agent, in response, learns slightly different policies: a more defensive policy in Capture, prioritizing the protection of its flag, and a more aggressive policy in Mix, focusing more on capturing the red flag.
+This policy difference is reflected in the weights in the constraint clause, where the $\neg\psi_{\mathrm{ra,bf}}$ predicate receives a higher weight in the Capture explanation (0.70) than in the Mix explanation (0.56), due to the Capture policy’s prioritization of flag protection.
+Conversely, the Mix explanation shows a higher weight for the $\psi_{\mathrm{ba,rf}}$ predicate (0.44) than the Capture explanation (0.30), due to its prioritization of flag capture.
 Note that $\psi_{\mathrm{ba,rf}}$ appears in both the task and constraint clauses in these explanations.
 This duplication results from our design choice to fix equal weights (0.5) on the task and constraint clauses, which forces TLNet to express prioritization between the task and constraint by including relevant task predicates in the constraint clause.
-Although one could alternatively include the task and constraints weights within the optimized parameters, we found that doing so introduced training instability.
 Now, consider the Capture and Capture 0 scenarios, which differ in the reward to kill the red agent (0.5 vs.
@@
 More broadly, this result also demonstrates the importance of inferring explanations from actual observed behaviors of RL agents rather than simply looking at their specified reward function.
-TABLE III: Ablation Study of TLNet in the CtF Capture Scenario.
-Metrics Full No Filter No Regularization No Pruning
-Conciseness 0.408 0.289 0.396 0.147
-Consistency 0.325 0.175 0.233 0.2
-Strictness 0.117 0.106 0.117 0.063
-Accuracy 0.951 0.947 0.954 0.776
+TABLE V: Ablation Study of TLNet in the CtF Capture Scenario.
+Features Concise. Consist. Strict. Accuracy
+[…]
+[…]
+[…]
+No Pruning: $_{N_{\mathrm{w}}=0}$ 0.15±0.02 0.20 0.06±0.00 0.78±0.05
+[…]
 IV-C3 Importance of TLNet Components
-Finally, the ablation study results shown in Table III show that filtering, regularization, and pruning are all important to the performance of TLNet.
-Removing predicate filtering notably reduced conciseness (by 27.2%) and consistency (by 46.1%).
-Removing regularization decreased conciseness (by 2.94%) and significanbly reduced consistency (by 28.3%) with marginal accuracy improvement.
-Finally, removing the pruning step nearly tripled conciseness and halved strictness.
+Finally, the ablation study results shown in Table V show that filtering, regularization, and pruning are all important to the performance of TLNet.
+Removing predicate filtering notably reduced conciseness and consistency, removing regularization decreased conciseness and significantly reduced consistency with no accuracy improvement, and removing pruning largely reduced all metrics.
+We also tested an argmin approach to sample negative trajectories, where rollouts used the lowest-probability action in the target policy, and saw no significant difference in results compared to random sampling.
 V Conclusion
-We introduce a neuro-symbolic framework, based on the TLNet architecture, for generating wSTL explanations of robot policies.
-Our approach generates easy-to-understand explanations through predicate filtering, loss regularization, and weight pruning.
-To assess explanation quality beyond classification metrics, we also propose explainability metrics—conciseness, consistency, and strictness.
-Experiments in three simulated robotic environments show that our method produces explanations that better capture these metrics while retaining high classification accuracy.
-By bridging neural network learning and formal logic-based explanations, our work advances explainability in complex robotic applications, with future directions including more expressive wSTL structures, human-in-the-loop refinement, and explainability-by-design policy optimization.
+We introduce TLNet, a neuro-symbolic framework for generating wSTL explanations of robot policies through predicate filtering, loss regularization, and weight pruning.
+We also propose novel explainability metrics—conciseness, consistency, and strictness—to assess explanation quality beyond classification accuracy.
+Experiments in three simulated robotic environments show our method outperforms baselines on these metrics while retaining high classification accuracy.
+Future directions include real-world deployment with online inference (which would require improving sample efficiency and handling predicate uncertainty from sensor noise and partial observability), learning task-constraint weights with a stabilizer, improving negative data generation, exploring post-hoc enforcement of the task-constraint format, and user studies.